### v3

# Energy-Based Model for Accurate Estimation of Shapley Values in Feature Attribution

Abstract
Shapley value is a widely used tool in explainable artificial intelligence (XAI), as it provides a principled way to attribute contributions of input features to model outputs. However, estimation of Shapley value requires capturing conditional dependencies among all feature combinations, which poses significant challenges in complex data environments. In this article, EmSHAP (Energy-based model for Shapley value estimation), an accurate and efficient Shapley value estimation method, is proposed to estimate the expectation of Shapley contribution function under the arbitrary subset of features given the rest. By utilizing the ability of energy-based model (EBM) to model complex distributions, EmSHAP provides an effective solution for estimating the required conditional probabilities. To further improve estimation accuracy, a GRU (Gated Recurrent Unit)-coupled partition function estimation method is introduced. The GRU network captures long-term dependencies with a lightweight parameterization and maps input features into a latent space to mitigate the influence of feature ordering. Additionally, a dynamic masking mechanism is incorporated to further enhance the robustness and accuracy by progressively increasing the masking rate. Theoretical analysis on the error bound as well as application to four case studies verified the higher accuracy and better scalability of EmSHAP in contrast to competitive methods.
Index Terms:
Explainable artificial intelligence, feature attribution, Shapley value, energy-based model.
I Introduction
Deep learning (DL) has been extensively applied in many complex supervised learning problems like medical diagnosis [1], industrial intelligence [2], and financial modeling [3]. DL models outperform conventional methods in terms of accuracy in many cases. However, since most DL models are black boxes and have complicated structures, interpretation of the results becomes difficult. This limitation poses challenges in identifying key features that influence the DL model and understanding the relationship between input features and outputs. This interpretability requirement leads to the development of an active research area of Explainable Artificial Intelligence (XAI) [4].
In XAI, the concept of model interpretability lacks a clear mathematical formulation. Nevertheless, there is a general consensus on the formulation proposed by Ref.[5]. According to Ref.[5], an explanation algorithm is considered trustworthy if it accurately reveals the underlying reasoning behind a model’s decisions. In this regard, Shapley value-based explanation becomes popular due to its rigid mathematical foundation and theoretical basis. Originated from the field of cooperative games [6] and reintroduced as an explanation framework [7], Shapley value effectively overcomes the limitations of black-box models through the implementation of fair allocation schemes. Although the utilization of Shapley values for explaining DL models may sometimes lead to non-intuitive feature attributions [8, 9, 10], the ideal mathematical properties make it possible to fairly measure the contribution of each input in the model. Hence, Shapley value-based interpretable approaches still remain a popular choice in XAI.
A notable problem for Shapley-based explanation is its high computation load. By definition, the computational complexity of Shapley value increases exponentially with the number of variables. In order to accelerate computation, a series of statistical-based methods have been developed that can be roughly divided into three categories: regression-based methods, sampling-based methods, and structure model-based methods. Regression-based methods employ weighted linear regression to accelerate the estimation of Shapley value, with examples including SHAP (SHapley Additive exPlanation) [11] and KernelSHAP [12]. Sampling-based methods utilize Monte Carlo sampling to approximate the true Shapley value. These methods involve Monte Carlo approximation using randomly sampled permutations [13] or sampling based on random permutations of local features [14]. On the other hand, structure model-based methods leverage specific model structure constraints to approximate the Shapley value. Examples of such methods include tree Shapley [15] and graph Shapley [16]. More recently, a generative model for Shapley value estimation based on variational autoencoder with arbitrary conditioning (VAEAC) [17] has been developed, which uses a single model for all feature combinations.
Despite the advancements made in calculation of Shapley value, several significant issues still remain. For example, regression-based methods can improve the estimation efficiency of Shapley value, however, the inherent assumption of independence among variables may negatively affect its accuracy. Sampling-based methods require a large quantity of samplings to achieve sufficient accuracy, and hence have high computation load. Whilst structure model-based approaches demonstrate improved efficiency, their application scenarios are limited to specific models. For VAEAC, it involves an additional amortization error by using the VAE model [18, 19]. Another notable research gap is that there still lacks rigid theoretical analysis on the error bounds of different Shapley value-based approaches, making it difficult to quantitatively evaluate their accuracies.
In order to increase estimation accuracy and provide theoretical guidance on the error bound, this article presents the EmSHAP (Energy-based model SHapley value estimation). Similar to Ref. [17], EmSHAP introduces a generative model during the estimation of Shapley contribution function. It relies on the energy-based model(EBM) to obtain the expectation of Shapley contribution function of the deep learning model under some arbitrary subset of features given the rest. The use of EBM is motivated by its ability to model complex, high-dimensional distributions without explicit normalization, which provides a flexible and efficient framework for estimating the conditional distributions required in Shapley value computation. Unlike original EBM, the EBM in EmSHAP is decomposed into an energy network to approximate the unnormalized conditional density and a GRU network for approximation of proposal distribution in the partition function. The GRU network has a simple structure and is capable of capturing long-term dependencies, and the adopted dynamic masking scheme improves the flexibility. The contributions of the paper can be summarized: i) Methodological refinements on the EBM framework by introducing the GRU network and the dynamic masking scheme, resulting in more accurate Shapley value estimation; ii) Theoretical contribution on the analysis of error bounds for several state-of-the-art methods, which confirmed the tighter error bound of EmSHAP; iii) Extensive application experiments are conducted to verify the accuracy and scalability of different Shapley value estimation methods.
II Relate work
Shapley value is a concept originated from cooperative game theory and provides a fair and reasonable method for quantifying feature attribution contributions. As a result, they have garnered significant attention in the field of XAI. However, the calculation of exact Shapley values is an NP-hard problem [20]. To address this computational complexity, researchers have developed various approximation methods. This section provides a brief review on the different kinds of approximation methods in literature, which can be summarized into sampling-based, regression-based, gradient-based, and generative model-based methods.
Sampling-based methods: Sampling-based Shapley value estimation is among the first to be studied and remains the most widely used. The core principle is to randomly sample subsets of features and calculate the average contribution of the features to model output. Variants of sampling-based methods include Monte Carlo sampling [21], Monte Carlo sampling with random arrangements [22, 7], and Owen-based sampling [23]. Most of them are model-agnostic and do not involve feature removal in Shapley value estimation, hence are popular in practical applications. However, sampling-based methods may suffer from high variance during the sampling procedures. The recently proposed SVARM [24] addresses this issue by controlling the estimated variance within a reasonable range.
Regression-based methods: Regression-based methods estimate Shapley values by using regression models to approximate the marginal contribution of each feature to the output. These methods approximate the original black-box model by constructing a simpler surrogate model. Among these methods, KernelSHAP [11] is the most widely used. It transforms the Shapley value estimation into a weighted least squares problem through the Shapley kernel, leading to significantly enhanced estimation efficiency. Based on KernelSHAP, Covert et al. [12] proposed an improved version that can reduce estimation bias, and Jethani et al. [25] developed a computational efficient method called FastSHAP. Additionally, the more complex variants like generalized linear models [26] and kernel regression [27] also belong to regression-based approaches. Compared to sampling-based methods, regression-based methods offer higher computational efficiency and lower estimation variance. However, since most regression methods compute the marginal contribution instead of real contribution of features, they may introduce bias in the estimates.
Gradient-based methods: Gradient-based methods estimate Shapley values by utilizing the gradient of the model’s output with respect to its input features. These methods compute the Shapley value by assessing the change in output resulting from the introduction of each feature increment. DeepSHAP [11] and its variant G-DeepSHAP [28] belong to this category, they are designed to address the computational cost of exact Shapley value estimation. However, DeepSHAP and G-DeepSHAP require access to the gradient of the features, hence are only applicable to differentiable models.
Generative model-based methods: Different from regression-based methods, generative model-based methods estimate the conditional probability of features in Shapley value estimation. For example, the on-manifold Shapley method [29] estimates the Shapley value within the manifold space and employs VAE to compute conditional probabilities across different feature subsets. VAEAC [17], a variant of this approach, applies various masking techniques to the training data to train a model capable of generating any conditional probability distribution. The problem of VAE based generative models is that the amortization error [18] in VAE may bring bias to the Shapley value estimation.
A more comprehensive literature review on Shapley-based feature attribution methods can be found in Ref. [30].
Table I summarizes typical Shapley value feature attribution methods, categorized by their approximation strategy and feature removal technique. It can be seen from Table I that each method has its own advantages, however, they all pose certain limitation in terms of estimation accuracy, which motivated us to undertake the current research.
TABLE I: Summary of typical Shapley value feature attribution methods
Methods Approximation strategy Removal approach Bias-free Pros Cons
ApproShapley[22] Sampling None Simple and general High computation cost
SamplingSHAP[7] Sampling None
SVARM[24] Sampling None
Multilinear extension[23] Sampling Marginal
KernelSHAP[11] Regression Marginal Computationally efficient Assume feature independence
unbiased-KernelSHAP[12] Regression Marginal
FastSHAP[25] Regression Marginal
Generalized additive model[26] Regression Marginal
RKHS-SHAP[27] Regression Marginal
DeepSHAP[11] Gradient Marginal Computationally efficient Model-dependent
G-DeepSHAP[28] Gradient Marginal
On-manifold Shapley[29] Generative model Conditional Flexible Amortization error
VAEAC[17] Generative model Conditional
III Preliminaries
In this section, preliminary knowledge about Shapley value-based explanation framework and energy-based model is presented.
III-A Shapley value
Shapley value is a tool from game theory designed to fairly allocate the contribution generated by a coalition game among its individual players [6]. Suppose a total of $d$ players participate in a game, and these players form a set of $D=\{1,\cdots,i,\cdots,|D|\}$, with $|D|$ being the number of all players. The contribution/revenue generated by these players in this game can be quantified as $v(D)$, with $v(\cdot)$ being a predefined contribution function. Similarly, the contribution generated by a player subset $S(S\subseteq D)$ can be quantified as $v(S)$. The Shapley value is defined as the marginal contribution of the player to the payoff of all possible player subsets. The definition of the Shapley value is as follows.
$\small\phi_{i}=\sum_{S\subseteq D\backslash\{i\}}\frac{|S|!(|D|-|S|-1)!}{|D|!}(v(S\cup\{i\})-v(S)).$ (1)
Here, $\phi_{i}$ is the Shapley value of the $i$-th player and $|S|$ is the number of players in subset $S$. The definition of Eq.(1) renders Shapley value the following favorable properties.
  • Efficiency: The sum of Shapley values for players from a subset is equal to the difference between the contribution function of the collection of these players and the contribution function of the empty set, i.e. $\sum_{i=1}^{|S|}\phi_{i}=v(S)-v(\varnothing)$;
  • Symmetry: For any coalition $S$ not containing $i_{1}$ and $i_{2}$, if player $i_{1}$ and player $i_{2}$ satisfy $v(S\cup\{i_{1}\})=v(S\cup\{i_{2}\})$, then $\phi_{i_{1}}=\phi_{i_{2}}$;
  • Dummy: Regardless of which coalition of players it is added to, a player $i$ does not change the contribution value, that is $v(S)=v(S\cup\{i\})$ for any coalition $S$, then $\phi_{i}=0$;
  • Additivity: For a game with combined contribution function $v_{1}+v_{2}$, the Shapley value can be calculated by $\phi_{i_{1}}+\phi_{i_{2}}$.
III-B Shapley value-based model explanation
The Shapley value, grounded in a rigorous mathematical framework and underpinned by fairness principles, serves as a powerful tool for feature attribution. By conceptualizing features as players and the model output as the value function, Shapley value offers a principled tool for quantifying the contribution of each feature to the model’s output, thereby enhancing interpretability of complex models.
Let $\tilde{f}:\mathbf{x}\in\mathbb{R}^{|D|}\rightarrow y\in\mathbb{R}$ be an unknown data generating function that generates a set of $n$ measurements $\mathbf{X}=\left\{\mathbf{x}^{t}\right\}^{n}_{t=1},\mathbf{x}^{t}\in\mathbb{R}^{|D|}$ of input features $\mathbf{x}$ and $\mathbf{Y}=\left\{y^{t}\right\}_{t=1}^{n},y^{t}\in\mathbb{R}$ of output feature $y$, and $y^{t}=\tilde{f}(\mathbf{x}^{t})$. $D$ is the input feature set and $|D|$ is the number of features. In a classification/predictive task, the purpose is to approximate the unknown function $\tilde{f}$ using a specific model $\hat{y}=f(\mathbf{x})$ to minimize the approximation error between $y$ and $\hat{y}$ as well as the out-of-sample generalization error. The approximation model $f(\cdot)$ can be of any type of machine learning models, in this paper, deep learning models are mainly considered.
In Shapley value-based model explanation framework, the $|D|$ input features are related to the $|D|$ players, the classification/predictive model $f(\mathbf{\cdot})$ relates to the cooperative game and the model output $\hat{y}$ can be regarded as the sum of contribution(or Shapley values) from all features in $\mathbf{x}$. Noting that the framework is model-agnostic, it can be used to interpret a wide variety of deep learning models.
The output for a specific sample $\mathbf{x}^{t}\in\mathbb{R}^{|D|\times 1}$ in $\mathbf{X}$ can be expressed as the sum of the Shapley values associated with each input feature, obtained by decomposing the model output. Specifically, this decomposition provides the following expression for the output corresponding to $\mathbf{x}^{t}$,
$\small f(\mathbf{x}^{t})=\phi_{0}+\sum^{|D|}_{i=1}\phi_{i}^{t}.$ (2)
Here, $\phi_{0}$ is the base value of the model outputs and can be regarded as the expectation of global output [11], with $\phi_{0}=\mathbb{E}[f(\mathbf{x})]$, $\phi_{i}^{t}$ is the Shapley value of the $i$-th input feature $x^{t}_{i}$. Eq.(2) highlights the disparity between the model output and the global output by employing the Shapley value associated with each feature in $\mathbf{x}^{t}$.
To calculate $\phi_{i}^{t}$, an appropriate contribution function $v(S)$ should be determined. Since $f(\mathbf{x}^{t})$ is the sum of contribution from all input features, we can easily relate $v(\cdot)$ to $f(\cdot)$, that is, $v(D)=f(\mathbf{x}^{t})$ for a specific sample and $\bar{v}(D)=\mathbb{E}[f(\mathbf{x})]$ for global output. Following Ref. [11], the contribution function can be reformed as the expected outcome of $f(\mathbf{x})$ conditioned on the variables in $S$ taking on specific values $\mathbf{x}_{S}^{t}$ as follows.
$v(S)$ $=\mathbb{E}[f(\mathbf{x})|\mathbf{x}_{S}=\mathbf{x}_{S}^{t}]=\mathbb{E}[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S})|\mathbf{x}_{S}=\mathbf{x}_{S}^{t}]$ (3)
[…]
where $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ is the conditional density of $\mathbf{x}_{\bar{S}}$ given $\mathbf{x}_{S}^{t}$, $\mathbf{x}_{S}$ and $\mathbf{x}_{S}^{t}$ contains the part of inputs $\mathbf{x}$ and $\mathbf{x}^{t}$ corresponding to features inside $S$, $\mathbf{x}_{\bar{S}}$ and $\mathbf{x}_{\bar{S}}^{t}$ relate to the part of inputs in $\mathbf{x}$ and $\mathbf{x}^{t}$ corresponding to features outside $S$, $S\cup\bar{S}=D$ and $S\cap\bar{S}=\varnothing$.
The Shapley value provides a fair assessment of each variable’s contribution to the model output. However, from Eq. (1) it can be seen that its estimation involves high computation load as there are a total of $2^{|D|-1}$ combinations of $S$. The computational complexity will increase exponentially as the number of input features grows.
III-C Energy-based model
The energy-based models (EBMs) are originated from statistical physics [31]. EBMs are a class of probabilistic model where the probability distribution of data is determined by an energy function. The primary advantage of EBMs is their universal approximation capability [32, 33]. This allows EBMs to model complex, high-dimensional, and multi-modal distributions, capturing intricate dependencies within the data. Unlike traditional parametric models, which rely on restrictive assumptions about the data(e.g., predefined distributions or fixed functional relationships), EBMs do not impose such constraints. As a result, they are well-suited for tasks involving complex or unknown data distributions. Assuming the input features are sampled from an unknown data distribution, i.e., $\mathbf{x}\sim p_{data}(\mathbf{x})$, the goal of energy-based model is to learn an unnormalized density $\exp(-g_{\theta}(\mathbf{x}))$ as a parametric approximation to $p_{data}(\mathbf{x})$.
$\small p_{data}(\mathbf{x})\approx p_{\theta}(\mathbf{x})=\frac{\exp(-g_{\theta}(\mathbf{x}))}{Z},$ (4)
[…]
The parameters in an energy-based model in Eq.(4) can be estimated by maximum likelihood estimation (MLE). Specifically, if $\mathbf{x}^{t}\sim p_{data}(\mathbf{x})$ for $t=1,2,...,n$, the likelihood function can be expressed as,
$\small\mathcal{L}(\theta)=\frac{1}{n}\sum_{t=1}^{n}\text{log}p_{\theta}(\mathbf{x}^{t})=\mathbb{E}_{\mathbf{x}\sim p_{data}(\mathbf{x})}[\text{log}p_{\theta}(\mathbf{x})].$ (5)
Since the energy function $g_{\theta}(\mathbf{x})$ in Eq.(4) can use different network architectures, it can approximate any kind of data distribution. In addition, the energy function can be multimodal without resorting to mixture distributions and does not differentiate between discrete and continuous distributions. These advantages make the energy-based model more flexible than traditional methods like kernel density estimation (KDE) [37].
IV Methodology
By definition of Shapley value in Eq.(1), it can be seen that the most important step is to estimate the contribution function $v(S)$, which can be reformulated as the conditional expectation in Eq.(3). The conditional expectation can be empirically approximated by techniques like Monte Carlo integration provided the conditional probability $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}^{t})$ is known.
[…]
$\approx\hat{v}(S)=\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t}),$
where $\mathbf{x}_{\bar{S}}^{(k)}\sim p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ for $k=1,2,...,K$, and $K$ is the number of Monte Carlo sampling. It is worth noting that estimation of conditional density $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ can be achieved by techniques like KDE. However, KDE is computationally challenging for high-dimensional inputs. More importantly, using KDE requires estimation of $2^{|D|-1}$ conditional densities as there are a total of $2^{|D|-1}$ combinations of $S$.
IV-A Network structure
To achieve an accurate estimation of conditional probability $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$, an energy-based model called EmSHAP is presented here, which is capable of estimating all conditional densities using a single model. It has the network structure as shown in Fig. 1.
Figure 1: Structure of GRU coupled energy-based model for estimating Shapley value
In Fig. 1, data batches of input features are fed into the energy network to estimate the energy function $\text{exp}(-g_{\theta}(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}))$, which can be used to approximate the conditional probability $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ after divided by the partition function. Here, the energy function $g_{\theta}$ can be designed as an unnormalized MLP with skip connections [38]. The partition function, on the other hand, needs to be approximated by Monte Carlo integration from the proposal distribution $q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$. The proposal distribution can be estimated using an autoregressive approach [39], however, its estimation accuracy will be negatively affected by feature orderings [^1] Feature ordering refers to the arrangement of input features in a specific sequence when fed into a neural network.. Instead, a GRU network with a dynamic masking scheme is used to estimate the proposal distribution, which outputs parameters related to the estimated proposal distribution and a context vector $\gamma$. The context vector is used to connect the energy network and the GRU network. With the conditional probability $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ obtained, the contribution function can be estimated from Eq.(6), and Shapley value can be subsequently obtained from Eq.(1). This network structure is under the framework of energy-based model, which can universally approximate any probability density according to Ref. [40]. Hence the proposed model structure inherits the good approximation ability of EBM. Comparing to the original EBM, a GRU network is used to generate the proposal distribution and integrate importance sampling to effectively estimate the partition function. The GRU sequentially generates the distribution of latent variables based on the masked variables and ultimately combines these into the proposal distribution. By capturing the intrinsic information associated with the sequential masking of variables, the GRU network ensures that the generated proposal distribution is more accurate, thereby enabling the model to better approximate complex conditional probability densities. The details of the network structure will be explained in Subsection IV-B.
IV-B GRU coupled energy-based model
When using the EBM to approximate the conditional density $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$, we are interested in estimating the conditional density of inputs in $\bar{S}$ conditioned on the remaining observed inputs (in $S$) for all possible $\bar{S}$. According to Eq.(4), $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ can be approximated using the EBM as follows.
$\small p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})\approx\frac{\exp(-g_{\theta}(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}))}{Z_{{\bar{S}};\mathbf{x}_{S}}}.$ (7)
The right hand side of Eq.(7) can be transformed to a probability once $Z_{\bar{S},\mathbf{x}_{S}}$ is known. However, it is always intractable to directly calculate $Z_{\bar{S},\mathbf{x}_{S}}$. One solution is to introduce a proposal conditional distribution $q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ that is close to $Z_{\bar{S},\mathbf{x}_{S}}$, then use Monte Carlo integration to sample from $q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ to approximate $Z_{\bar{S},\mathbf{x}_{S}}$.
[…]
$\approx\frac{1}{\tilde{K}}\sum_{\tilde{k}=1}^{\tilde{K}}\frac{\text{exp}(-g_{\theta}(\mathbf{x}_{\bar{S}}^{(\tilde{k})},\mathbf{x}_{S}))}{q(\mathbf{x}_{\bar{S}}^{(\tilde{k})}|\mathbf{x}_{S})},\mathbf{x}_{\bar{S}}^{(\tilde{k})}\sim q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}).$
Here $\hat{Z}_{{\bar{S}};\mathbf{x}_{S}}$ is the approximate partition function, $\tilde{K}$ is the number of Monte-Carlo sampling. The above procedure is also called importance sampling. When $q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ and $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ are close, such a sampling estimation can be regarded as unbiased. Thus, an appropriate proposal distribution $q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ should be defined, which is desired to be close to $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ as much as possible and have a parametric expression that is easy to be inferred.
The proposal distribution $q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ can be obtained through an autoregressive approach [39],
$\small q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})\approx\prod_{l=1}^{|{\bar{S}}|}q(x_{{\bar{S}_{l}}}|\mathbf{x}_{S\cup\bar{S}_{<l}}).$ (9)
Eq.(9) uses the chain rule of conditional probability to obtain the proposal distribution, so that estimation of the proposal distribution can be converted into the multiplication of a series of one-dimensional conditional densities $q(x_{{\bar{S}}_{l}}|\mathbf{x}_{S\cup\bar{S}_{<l}})$, which is easy and tractable.
An important issue arises here, namely, the feature ordering of the conditional probabilities $q(x_{{\bar{S}}_{l}}|\mathbf{x}_{S\cup\bar{S}_{<l}})$ will have a significant impact on the estimation results. For example, for a predictive task with three inputs $x_{1}$, $x_{2}$ and $x_{3}$, when calculating the conditional density $q(x_{2},x_{3}|x_{1})$, it can be obtained as $q(x_{2},x_{3}|x_{1})=q(x_{3}|x_{1},x_{2})q(x_{2}|x_{1})$ or $q(x_{2},x_{3}|x_{1})=q(x_{2}|x_{1},x_{3})q(x_{3}|x_{1})$ according to different feature orderings. Theoretically, $q(x_{3}|x_{1},x_{2})q(x_{2}|x_{1})$ is equal to $q(x_{2}|x_{1},x_{3})q(x_{3}|x_{1})$, however, since we use the energy-based model to approximate these distributions, the two orderings may produce different results. This is problematic if the number of input features is large. Though Ryan et al. [39] used plenty of random training to alleviate this problem, the solution is still unsatisfactory.
In order to better estimate the conditional distribution and subsequently the Shapley value, a GRU network is applied here. Let $|{\bar{S}}|$ denote the number of masked variables in $\mathbf{x}_{\bar{S}}$ and $x_{\bar{S}_{j}}$ be the $j$-th($j=1,\dots,|{\bar{S}}|$) masked variable, the conditional density $q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ can be approximated by multiplying the density functions of $|{\bar{S}}|$ Gaussian distributed auxiliary variables $\mathbf{z}_{j}$, whose mean vector and covariance matrix can be generated through the GRU network in Fig. 2.
Figure 2: GRU network for proposal distribution estimation
In Fig. 2, the unmasked features $\mathbf{x}_{S}$ together with the $j$-th masked feature $x_{\bar{S}_{j}}$ are fed into the GRU cell to generate the $j$-th hidden state $\mathbf{h}_{j}$ based on the information of the $j-1$-th hidden state $\mathbf{h}_{j-1}$, which can be summarized as follows.
$\mathbf{h}_{j}=\text{GRU}\left(\left[\mathbf{x}_{S},x_{\bar{S}_{j}}\right],\mathbf{h}_{j-1}\right).$ (10)
It should be noted that a total of $|D|-1$ zeros are combined with $x_{\bar{S}_{j}}$ in Fig. 2 to align with the dimension of $\mathbf{x}_{S}$. The hidden state has a dimension of $2|D|+|\bm{\gamma}|$, with the first $|D|$ elements forming the mean vector of the auxiliary variables $\mathbf{z}_{j}$, the second $|D|$ elements forming the diagonal vector of $\mathbf{z}_{j}$’s covariance matrix, and the last $|\bm{\gamma}|$ elements forming the context vector $\bm{\gamma}_{j}$. The hidden state $\mathbf{h}_{j}$ can be denoted as follows.
$\left[\bm{\mu}_{\mathbf{z}_{j}}^{T},\bm{\sigma}_{\mathbf{z}_{j}}^{T},\bm{\gamma}_{j}^{T}\right]^{T}=\mathbf{h}_{j}.$ (11)
Here $\bm{\mu}_{\mathbf{z}_{j}}$ is the mean vector, $\bm{\sigma}_{\mathbf{z}_{j}}$ is the vector containing the diagonal elements of $\mathbf{z}_{j}$’s covariance matrix. The context vector $\bm{\gamma}_{j}$ is used to connect the GRU network with the energy network.
The auxiliary variables $\mathbf{z}_{j}$ has a Gaussian distribution as follows.
$q(\mathbf{z}_{j})\sim\mathcal{N}(\mathbf{z}_{j}|\bm{\mu}_{\mathbf{z}_{j}},\bm{\Sigma}_{\mathbf{z}_{j}}).$ (12)
Here $\bm{\Sigma}_{\mathbf{z}_{j}}$ is the diagonal covariance matrix with diagonal elements being $\bm{\sigma}_{\mathbf{z}_{j}}$.
The masked feature $x_{\bar{S}_{j}}$ is fed into the GRU cell one by one to yield a series of auxiliary variable $\mathbf{z}_{j}$, with its distribution generated by $\mathbf{h}_{j}$. Benefiting from the inherent gating mechanism of the GRU network, $\mathbf{h}_{j}$ consistently incorporates information from both previously masked features and the current masked feature. Consequently, the product of the densities of all generated auxiliary variables can be used to approximate the desired conditional probability distribution.
$q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})\approx\prod_{j=1}^{|\bar{S}|}q(\mathbf{z}_{j}).$ (13)
Throughout the training process, whenever a masked feature is fed into the GRU network, the corresponding information is captured and stored in the hidden state $\mathbf{h}$. The hidden state $\mathbf{h}$, on the other hand, can be globally optimized by minimizing the loss function, making it insensitive to the orderings of feature masking. Hence, compared to Eq.(9), the introduction of the hidden state $\mathbf{h}$ allows GRU to better capture global dependencies among features, which significantly reduces the impact of feature ordering.
Another notable issue here is that Gaussian distribution is used as the prior distribution of the auxiliary variables. Although it is not a perfect prior for complex data or model, it is widely used in importance sampling. According to Ref. [41], importance sampling can effectively approximate the partition function as long as the proposal distribution has similar shape or overlapping region with the real distribution, and Gaussian distribution well fits this condition.
With the proposal distribution $q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ determined, the loss function of the GRU network can be established according to the maximum likelihood criterion.
$\quad\mathcal{J}({\Theta};\mathbf{x})=-\mathcal{L}({\Theta};\mathbf{x})$ (14)
$=-\frac{1}{\tilde{K}}\sum_{\tilde{k}=1}^{\tilde{K}}\log p(\mathbf{x}^{(\tilde{k})}_{\bar{S}}|\mathbf{x}_{S},{\Theta})-\text{log}q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S},{\Theta}).$
Here ${\Theta}$ is the parameter set.
Using the GRU network structure in Fig. 2, the conditional densities for any combination of $\mathbf{x}_{\bar{S}}$ given $\mathbf{x}_{S}$ can be obtained so that a single model can estimate all the conditional densities and repetitive operation of conditional density estimations can be avoided.
In addition, the difference in the variance of the data sampled from the proposal conditional distribution and the true conditional distribution has an explicit expression. Specifically, the data sampled from the true conditional distribution has the following variance,
$\small\text{Var}_{\mathbf{x}_{\bar{S}}\sim p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}[Z_{{\bar{S}};\mathbf{x}_{S}}]=\mathbb{E}_{\mathbf{x}_{\bar{S}}\sim p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}[Z_{{\bar{S}};\mathbf{x}_{S}}^{2}]-\left(\mathbb{E}_{\mathbf{x}_{\bar{S}}\sim p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}[Z_{{\bar{S}};\mathbf{x}_{S}}]\right)^{2}.$ (15)
Meanwhile, the variance of the data sampled from the proposal conditional distribution is,
$\quad\text{Var}_{\mathbf{x}_{\bar{S}}\sim q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}\left[Z_{{\bar{S}};\mathbf{x}_{S}}\frac{p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}{q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}\right]$ (16)
$=\mathbb{E}_{\mathbf{x}_{\bar{S}}\sim q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}\left[\left(Z_{{\bar{S}};\mathbf{x}_{S}}\frac{p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}{q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}\right)^{2}\right]$
$\quad-\left(\mathbb{E}_{\mathbf{x}_{\bar{S}}\sim q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}\left[Z_{{\bar{S}};\mathbf{x}_{S}}\frac{p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}{q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}\right]\right)^{2}$
$=\mathbb{E}_{\mathbf{x}_{\bar{S}}\sim p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}\left[Z_{{\bar{S}};\mathbf{x}_{S}}^{2}\frac{p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}{q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}\right]-\left(\mathbb{E}_{\mathbf{x}_{\bar{S}}\sim p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}[Z_{{\bar{S}};\mathbf{x}_{S}}]\right)^{2}.$
Subtract Eq.(15) and Eq.(16), the following equation is obtained:
$\quad\text{Var}_{\mathbf{x}_{\bar{S}}\sim p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}[Z_{{\bar{S}};\mathbf{x}_{S}}]-\text{Var}_{\mathbf{x}_{\bar{S}}\sim q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}[Z_{{\bar{S}};\mathbf{x}_{S}}]$ (17)
$=\mathbb{E}_{\mathbf{x}_{\bar{S}}\sim p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}\left[Z_{{\bar{S}};\mathbf{x}_{S}}^{2}\right]-\mathbb{E}_{\mathbf{x}_{\bar{S}}\sim p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}\left[Z_{{\bar{S}};\mathbf{x}_{S}}^{2}\frac{p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}{q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}\right].$
As can be seen from Eq.(17), when the proposal distribution $q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ is close to the true distribution $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$, the discrepancy in variance between the approximate partition function $Z_{{\bar{S}};\mathbf{x}_{S}}$ and the true partition function approaches zero.
IV-C Dynamic masking scheme
Calculation of Shapley values requires estimating the conditional expectations of contribution function under different feature combinations. This involves repeatedly evaluating the impact on the model output by including or excluding specific input features. The masking scheme provides an effective way to perform this evaluating procedure. In the masking scheme, a Bernoulli distribution is used to determine whether an input feature is included or excluded, which can be described as follows.
$\begin{gathered}b_{i}\sim Bernoulli(\zeta),\\ \tilde{x}_{i}=x_{i}\cdot(1-b_{i}).\end{gathered}$ (18)
Here, $\zeta$ is the masking probability, and $b_{i}$ is a binary value generated from the Bernoulli distribution, with $b_{i}=0$ indicating $x_{i}$ being included in the feature set $S$, otherwise $x_{i}$ is masked and excluded from the feature set $S$. The masking probability $\zeta$ is an important parameter that needs to be manually tuned, for example, $\zeta$ can be set as 0.5 by trial and error. However, using a fixed masking probability is problematic, as it may not generalize to a wide range of potential feature combinations, especially when the number of possible feature combinations is very large. Here, a dynamic masking scheme is introduced by gradually increasing $\zeta$, for example, from 0.2 to 0.8, to cover a wider range of feature combinations. Another benefit of using the dynamic masking scheme is that there is no need for manually tuning the optimal masking rate.
In the dynamic masking scheme, the $\zeta$ value is linearly increased during the training epochs of the model. This ensures EmSHAP to focus on the simpler scenarios with fewer features being masked in the initial stage, and progressively transiting to more complex scenarios with a high proportion of features being masked. This process can be formulated as follows.
$\small\zeta_{e+1}=\zeta_{e}+\Delta.$ (19)
Here $e$ is the epoch index in model training and $\Delta$ is the increment of $\zeta$ at each epoch. As the training epoch grows larger, the masking rate increases linearly from a preset initial value of $\zeta_{min}$ until it reaches the preset maximum masking rate $\zeta_{max}$.
By applying the dynamic masking scheme, the GRU network recursively processes the masked input features to capture dependencies among variables. At each step, GRU updates its hidden state by incorporating historical masking information and generates the conditional probability estimate for the current step. This process enables GRU to leverage its internal gating mechanisms to retain and propagate prior masking information, thus effectively capturing feature dependencies without relying on a fixed feature order. The detailed procedures for training EmSHAP are listed in Algorithm 1.
Algorithm 1 Training procedures of EmSHAP
0:  Training data $\mathbf{x}$, training epochs $\mathcal{E}$
  Initialize: Energy network parameters, GRU network parameters, and dynamic masking scheme, $e=1$.
  while $e\leq\mathcal{E}$ do
     $[\mathbf{x}_{\bar{S}},\mathbf{x}_{S}]=$ dynamic masking$(\mathbf{x})$
     for $j=1$ to $|\bar{S}|$ do
        $\mathbf{h}_{j}=\text{GRU}\left(\left[\mathbf{x}_{S},x_{\bar{S}_{j}}\right],\mathbf{h}_{j-1}\right)$
        $\left[\bm{\mu}_{\mathbf{z}_{j}}^{T},\bm{\sigma}_{\mathbf{z}_{j}}^{T},\bm{\gamma}_{j}^{T}\right]^{T}=\mathbf{h}_{j}$
        $q(\mathbf{z}_{j})\sim\mathcal{N}(\mathbf{z}_{j}|\bm{\mu}_{\mathbf{z}_{j}},\bm{\Sigma}_{\mathbf{z}_{j}})$
     end for
     $q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})=\prod_{j=1}^{|\bar{S}|}q(\mathbf{z}_{j})$
     Sample from $q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ to get $\mathbf{x}_{\bar{S}}^{(\tilde{k})}$
     energy network$([\mathbf{x}_{\bar{S}}^{(\tilde{k})},\mathbf{x}_{S}],\bm{\gamma}_{|\bar{S}|})$rightarrow $\exp(-g_{\theta}(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}))$
     Calculate partition function using Eq.(8)
     Calculate loss using Eq.(14)
     $e=e+1$
  end while
V Analysis on Error bound
In this section, the error bound of the contribution function in the proposed EmSHAP is discussed. Theorem 1 indicates that the expectation of the mean absolute deviation (MAD) between the estimated contribution function $\hat{v}(S)$ and the true contribution function $v(S)$ has an upper bound of $\frac{\sqrt{\pi}}{\sqrt{2K}}+\epsilon_{2}$, where $K$ is the number of Monte Carlo sampling, $\epsilon_{2}$ is a small positive number related to the absolute probability density ratio between the estimated conditional density and the true data-generating density. For comparison, Theorems 2 and 3 presents the error bounds of two competitive methods, i.e., KernelSHAP [12] and VAEAC [17].
To begin with, the following assumptions are introduced to facilitate the analysis.
Assumption 1.
The parameters of the estimated conditional density $p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ are unbiased estimates of the parameters of the true data-generating density $p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$, that is, for any small positive $\epsilon_{1}$, the probability density ratio can be described as follows,
$\small\left|\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-1\right|\leq\epsilon_{1}.$ (20)
Assumption 2.
The model $f(\mathbf{x})$ is continuous and second-order differentiable. Once $\mathbf{x}_{\bar{S}}$ estimated from EmSHAP is close to the real value $\mathbf{x}_{\bar{S}}^{t}$, the Taylor expansion of $f(\mathbf{x}_{\bar{S}},\mathbf{x}_{s}^{t})$ on $f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{s}^{t})$ can be expressed as follows.
$f(\mathbf{x}_{\bar{S}},\mathbf{x}_{s}^{t})=f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{s}^{t})+f^{\prime}(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{s}^{t})(\mathbf{x}_{\bar{S}}-\mathbf{x}_{\bar{S}}^{t})+o(\mathbf{x}_{\bar{S}}-\mathbf{x}_{\bar{S}}^{t}),$ (21)
where $o(\mathbf{x}_{\bar{S}}-\mathbf{x}_{\bar{S}}^{t})\leq\epsilon_{1}^{\prime}$, $\epsilon_{1}^{\prime}$ is a small positive number.
Assumption 3.
The contribution value $v(S)$ of any subset $S$ and the output value $\hat{y}=f(\mathbf{x})$ are bounded within $[0,1]$.
Assumption 1 holds as EBM can be flexibly parameterized for arbitrary probability distributions [42, 43]. More specifically, the proposed GRU coupled energy-based model consists of an energy network and a GRU network. The energy network is an MLP with a skip connection and the GRU network is a type of gated recurrent network. According to Theorem 2.5 in Ref.[44], with weighted path norm uniformly bounded, by setting appropriate network width and depth, MLP with skip connection can approximate any function in Barron space [^2] Barron space is a function space that consists of functions which can be approximated by smooth combinations of simple basis functions.. On the other hand, Theorem 3.2 in Ref.[45] shows the universal approximation ability of GRU.
Before analyzing the upper error bound of EmSHAP, we first demonstrate the asymptotic convergence of $\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]$ and $\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]$, followed by the derivation of their error bounds.
V-A Asymptotic convergence of $\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]$ and $\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]$
Proposition 1.
(Consistency) From the weak law of large numbers, when the sampling time $K$ is large enough, $\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})$ is convergent by probability to the conditional expectation $\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]$, that is, for any positive $\epsilon_{2}$, we have,
$\mathop{\lim}_{K\rightarrow+\infty}\mathbb{P}\left(\left|\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|\leq\epsilon_{2}\right)=1.$ (22)
Proof.
Let $\bar{f}(\mathbf{x}^{(k)}_{\bar{S}},\mathbf{x}_{S}^{t})=\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})$ and $\text{Var}(f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t}))=\sigma$ for all $k$, the variance of $\bar{f}(\mathbf{x}^{(k)}_{\bar{S}},\mathbf{x}_{S}^{t})$ is as follows.
$\quad\text{Var}\left(\bar{f}(\mathbf{x}^{(k)}_{\bar{S}},\mathbf{x}_{S}^{t})\right)=\text{Var}\left(\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})\right)$ (23)
$=\frac{1}{K^{2}}\text{Var}\left(\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})\right)=\frac{K\sigma^{2}}{K^{2}}=\frac{\sigma^{2}}{K}.$
By using Chebyshev’s inequality [46] on $\bar{f}(\mathbf{x}^{(k)}_{\bar{S}},\mathbf{x}_{S}^{t})$, we have,
$\small\mathbb{P}\left(\left|\bar{f}(\mathbf{x}^{(k)}_{\bar{S}},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right|\leq\epsilon_{2}\right)\geq 1-\frac{\sigma^{2}}{K\epsilon_{2}^{2}}.$ (24)
As $K$ approaches infinity, Eq.(22) can be obtained. ∎
Proposition 2.
The expectation of estimated conditional density $p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ is convergent by probability to the expectation of true data-generating density $p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$, that is, for any positive $\epsilon_{2}$, we have,
$\mathbb{P}$ $\left(\left|\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|\leq\epsilon_{2}\right)=1.$ (25)
Proof.
To prove Proposition 2, Taylor expansion for the expectation function is used. The absolute value of the difference between the estimated conditional expectation and the true conditional expectation is as follows.
$\quad\left|\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|$ (26)
$=\left|\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\left(f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}^{t}_{S})+f^{\prime}(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}^{t}_{S})(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})+o(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})\right)\right]\right.$
$\quad\left.-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|$
$=\left|\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}^{t}_{S})\right]-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right.$
$\quad\left.+\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f^{\prime}(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}^{t}_{S})(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})+o(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})\right]\right|.$
According to Assumption 1, the first term of the last equation in Eq.(26) is bounded by $\epsilon_{1}$, that is, $\left|\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}^{t}_{S})\right]-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|\leq\epsilon_{1}$. In addition, since the parameters of $p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ are unbiased estimates of $p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$, we have $\mathbb{E}\left[\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}}\right]=0$ and $\mathbb{E}\left[o(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})\right]\leq\epsilon_{1}^{\prime}$. Thus, for the given smalls positive $\epsilon_{1}$ and $\epsilon_{1}^{\prime}$, there always exists a small positive $\epsilon_{2}$ and Eq.(25) can be obtained. ∎
Proposition 1 and Proposition 2 imply the convergence of Monte Carlo sampling and the proposed EmSHAP. Furthermore, the following corollary can be obtained.
Corollary 1.
When the sampling number $K$ goes to infinity, the mean of the samplings converges to the true conditional expectation, for a small positive number $\epsilon_{2}$, we have,
$\small\mathop{\lim}_{k\rightarrow\infty}\mathbb{P}\left(\left|\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|\leq\epsilon_{2}\right)=1.$ (27)
The larger $K$ is, the closer the mean of samplings is to the true conditional expectation, which indicates the lower bound of the difference between the estimation conditional expectation and the true conditional function is zero.
V-B Error bound of approximating $\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]$ and $\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]$
Here the error bound of approximating $\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]$ is first discussed.
Lemma 1.
(Hoeffding’s inequality) [47] Let $u_{1}$,…$u_{K}$ be i.i.d. random variables such that $a\leq u_{k}\leq b$. Then for any $\epsilon_{3}>0$,
$\small\mathbb{P}\left(\left|\frac{1}{K}\sum_{k=1}^{K}\left(u_{k}-\mathbb{E}[u_{k}]\right)\right|\geq\epsilon_{3}\right)\leq 2\exp\left(-\frac{2K\epsilon_{3}^{2}}{(b-a)^{2}}\right).$ (28)
Hoeffding’s inequality is an important technique for establishing upper bounds on the probability of sums involving bounded random variables deviating significantly from their expected values. By using Lemma 1, the following corollary is obtained.
Corollary 2.
(corollary of Hoeffding’s inequality) For any $\epsilon_{3}>0$, we have,
$\mathbb{P}\left(\left|\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|\geq\epsilon_{3}\right)$ (29)
$\leq 2\exp\left(-2K\epsilon_{3}^{2}\right).$
Proof.
Since the output value is bounded within $[0,1]$ according to Assumption 3, by applying Hoeffdings inequality, we can get,
$\quad\mathbb{P}\left(\left|\frac{1}{K}\sum_{k=1}^{K}\left(f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}[f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})]\right)\right|\geq\epsilon_{3}\right)$ (30)
$=\mathbb{P}\left(\left|\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}[f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})]\right|\geq\epsilon_{3}\right)$
$\leq 2\exp\left(-2K\epsilon_{3}^{2}\right).$
$\mathbf{x}^{(k)}_{\bar{S}}$ is sampled from $p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ and the expectation of $f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})$ is equal to $\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]$ for any sample. Thus, Eq.(29) is obtained. ∎
The error bound of approximating $\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]$ is discussed as follows.
Proposition 3.
For any small positive $\epsilon_{2}$, the following inequality can be constructed,
$\small\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\mathbb{E}\left|f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right|\right]\leq\epsilon_{2}.$ (31)
Proof.
The proof of Proposition 3 relies on the Taylor expansion as well, which is described as follows.
$\quad\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\left|f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{S}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right|\right]$
$=\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\left|\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})+f^{\prime}(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})\right.\right.\right.$
$\quad\left.\left.\left.+o(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})\right]-f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right|\right]$
$=\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\left|\left(\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-1\right)f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right.\right.$
$\quad\left.\left.+\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left(f^{\prime}(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})+o(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})\right)\right|\right]$
$=\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\left|\left(\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-1\right)f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right.\right.$
$\quad\left.\left.+o(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\right|\right]$
$\leq\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\left|\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-1\right|\cdot\left|f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right|\right.$
$\quad\left.+\left|o(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\right|\right]$
$\leq\epsilon_{1}f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})+\epsilon_{1}^{\prime}$ (32)
$\leq\epsilon_{2}.$
According to Assumptions 1 and 2, $\mathbb{E}\left[\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}}\right]=0$ and $\mathbb{E}\left[o(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})\right]\leq\epsilon_{1}^{\prime}$, we have,
$\small\frac{p_{\theta}(\mathbf{x}_{S}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{S}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left(f^{\prime}(\mathbf{x}^{t}_{S},\mathbf{x}_{S}^{t})(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})+o(\mathbf{x}_{\bar{S}}-\mathbf{x}^{t}_{\bar{S}})\right)\leq\epsilon_{1}^{\prime}.$ (33)
Then, according to Assumption 3, $f(\mathbf{x}^{t}_{S},\mathbf{x}_{S}^{t})$ is bounded within $[0,1]$, and Eq.(31) can be obtained. ∎
V-C Theoretical error bounds of EmSHAP and competitive methods
Based on the propositions and corollaries discussed above, the upper bound between the expectations of the mean absolute deviation for the estimation contribution function $\hat{v}(S)$ and the true contribution function $v(S)$ can be calculated as follows.
Theorem 1.
(EmSHAP estimation error upper bound) The expectation of the mean absolute deviation (MAD) between the estimation contribution function $\hat{v}(S)$ and the true contribution function $v(S)$ is upper bounded by $\frac{\sqrt{\pi}}{\sqrt{2K}}+\epsilon_{2}$, with $K$ being the number of sampling and $\epsilon_{2}$ a small positive number.
Proof.
$\mathbb{E}[|\hat{v}(S)-v(S)|]$ can be decomposed into two parts, namely the statistical error and the approximation error. The detailed derivation is given as follows.
$\quad\mathbb{E}[|\hat{v}(S)-v(S)|]$ (34)
$=\mathbb{E}\left[\left|\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\right]\right|\right]$
$=\mathbb{E}\left[\left|\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right.\right.$
$\quad\left.\left.+\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\right]\right|\right]$
$\leq\mathbb{E}\left[\left|\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|\right]$
$\quad+\mathbb{E}\left[\left|\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\right]\right|\right],$
where $\mathbf{x}_{\bar{S}}\sim p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$. The first term in the last inequality of Eq.(34) is the statistical error and the second is the approximation error. The upper bound of the statistical error is first considered.
$\quad\mathbb{E}\left[\left|\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|\right]$ (35)
$=2\mathbb{E}\left[\left(\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right)_{+}\right]$
[…]
$=\frac{\sqrt{\pi}}{\sqrt{2K}},$
[…]
$\quad\mathbb{E}\left[\left|\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\right]\right|\right]$ (36)
$=\mathbb{E}\left[\left|\int f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}\right.\right.$
$\quad\left.\left.-\int f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}\right|\right]$
$=\mathbb{E}\left[\left|\int f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}\right.\right.$
$\left.\left.\quad-\int f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}\right|\right]$
$=\mathbb{E}\left[\left|\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\left(f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\right)\right]\right|\right]$
$\leq\mathbb{E}\left[\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\left|f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\right|\right]\right]$
$=\epsilon_{2}.$
By combining Eq.(35) and Eq.(36), the upper bound of $\mathbb{E}[|\hat{v}(S)-v(S)|]$ can be obtained, which is $\frac{\sqrt{\pi}}{\sqrt{2K}}+\epsilon_{2}$. ∎
We now discuss the tightness of the upper bound. The first inequality used is the triangle inequality for absolute values in Eq.(34). If the two functions on the right side of the inequality share the same sign (both positive or negative), that is,
$\small\begin{gathered}\left(\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right)\cdot\\ \left(\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\right]\right)\geq 0,\end{gathered}$ (37)
the bound derived from applying the absolute value triangle inequality is tight. Another inequality is the expectation inequality of absolute values in Eq.(36). This inequality shows that when the variable always has the same sign, the following holds.
$\small\begin{gathered}\sum_{\mathbf{x}_{\bar{S}}}f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\\ =\sum_{\mathbf{x}_{\bar{S}}}\left|f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\right|.\end{gathered}$ (38)
Hence, if $\mathbf{x}_{\bar{S}}$ satisfies both Eqs.(37) and (38), the upper bound is tight.
Theorem 2.
(KernelSHAP estimation error upper bound) For any fixed $|D|$ and when the sampling number $K$ goes to positive infinity, under the probability greater than $1-2((|D|+1)^{2}+(|D|+1))\exp(-2K\epsilon_{3}^{2})$, the expectation of the mean absolute deviation (MAD) between the estimation contribution function $\hat{v}(S)$ estimated by KernelSHAP [12] and the true contribution function $v(S)$ is upper bounded by,
$\mathbb{E}[|\hat{v}(S)-v(S)|]$ $\leq 2(|D|+1)(|S|+1)\frac{\sqrt{\pi}}{\sqrt{2K}}||\bm{\Sigma}^{*^{-1}}||_{F}$ (39)
$\quad+2(|D|+1)^{\frac{5}{2}}(|S|+1)\frac{\sqrt{\pi}}{\sqrt{2K}}||\bm{\Sigma}^{*^{-1}}||^{2}_{F}+\epsilon_{3}^{2},$
where $\epsilon_{3}$ is a small positive number. $\bm{\Sigma}^{*}\in\mathbb{R}^{K\times K}$ is the expectation of Shapley kernel symmetric matrix, more detailed information about this matrix can be found in Ref. [12].
Proof.
The detailed proof can be found in Appendix. A. ∎
Theorem 3.
(VAEAC estimation error upper bound) The expectation of the mean absolute deviation (MAD) between the contribution function $\hat{v}(S)$ estimated by VAEAC [17] and the true contribution function $v(S)$ is upper bounded by $\frac{\sqrt{\pi}}{\sqrt{2K}}+\epsilon_{2}+\delta$, where $K$ is the sampling number, $\epsilon_{2}$ is the probability density ratio between the estimation conditional probability $p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ and the optimal conditional probability $p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ estimated using VAE, $\delta$ is a positive number related to the amortization error of VAE.
$\small\left|\frac{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-1\right|\leq\delta.$ (40)
Proof.
The detailed proof can be found in Appendix. B. ∎
Theorems 2 and 3 analyzed the error bounds of two widely applied methods of KernelSHAP [12] and VAEAC [17]. Theorem 2 shows that as the sampling time $K$ increases, the first term of the right hand side of Eq.(39) vanishes, hence its error bound is governed by the second term. Since $||\bm{\Sigma}^{*^{-1}}||^{2}_{F}$ is a constant, it can be seen that the error bound of KernelSHAP is highly related to the dimension of input features $D$. That is to say, the estimation accuracy will decrease when the input feature dimension increases. On the other hand, comparing Theorem 1 and Theorem 3, it can be seen that the error bound of VAEAC is greater than the energy-based model as $\delta$ is always greater than zero. Hence it can be concluded that the proposed EmSHAP has a tighter error bound than KernelSHAP and VAEAC.
V-D Validation of Error Bound through a toy example
In order to give a more intuitive exhibition of the results in Theorem 1, a toy example is used to visualize the error bound 1.
Consider a simple example involving three inputs $\mathbf{x}=[x_{1},x_{2},x_{3}]$, with each feature $x_{i}(i=1,2,3)$ generated by a periodic function as follows.
$x_{i}(t)=\sin(it\pi)+\eta_{1}(t).$ (41)
Here, $t$ is the time instance ranging from 0 to 10, and is uniformly sampled into 5,000 points, $\eta_{1}$ is Gaussian distributed noise with mean 0 and variance 0.001. The output $y$ is obtained by a linear combination of the features of the input vector $\mathbf{x}$ with a weight vector $\mathbf{w}=[-0.4,-1.2,0.8]$, as follows.
$y(t)=\mathbf{w}^{T}\mathbf{x}(t)=-0.4x_{1}(t)-1.2x_{2}(t)+0.8x_{3}(t)+\eta_{2}(t),$ (42)
where $\eta_{2}$ is Gaussian noise with mean 0 and variance 0.001.
The first 80% of data samples are used to train a random forest model and EmSHAP, and the remaining 20% of the samples are used to verify Theorem 1. In this toy example, the statistical error is the difference between the predictions obtained using the mean of $K$ samples drawn from the EmSHAP output distribution and its expectation. The approximation error $\epsilon_{2}$ is the difference between the predictions derived from the expectation of the EmSHAP output distribution and the real data.
Fig 3 shows the curve of the contribution function obtained through EmSHAP under different feature combinations. From Fig 3, it can be seen that the approximate error is fixed for each feature combination, and the statistical error continues to decrease with the increase in the number of samplings. And the sum of the approximate and statistical errors is always below the error boundary. This validates the correctness of Theorem 1.
Figure 3: Toy example for validating error bound of EmSHAP
VI Case studies
In this section, feature attribution experiments are performed on four real-world tasks: image classification on MNIST [^3] MNIST dataset was acquired at http://yann.lecun.com/exdb/mnist/., question classification on TREC [^4] TREC dataset was acquired at https://cogcomp.seas.upenn.edu/Data/QA/QC/., tabular classification on adults’ income(ADTI) [^5] ADTI was acquired at https://archive.ics.uci.edu/dataset/2/adult., and time-series prediction of electric transformer’s oil temperature(ETT) [^6] ETT dataset was acquired at https://github.com/zhouhaoyi/ETDataset.. For the four tasks, EmSHAP is used to estimate Shapley values for feature attribution, which is compared with five competitive methods:
  • SamplingSHAP [7]: A Shapley value estimation method based on sampling from random permutation of features.
  • SVARM [24]: A sampling-based Shapley value calculation method without marginal contribution calculation.
  • KernelSHAP [12]: An effective method for calculating Shapley values which uses weighted least squares to calculate marginal Shapley values.
  • FastSHAP [25]: A fast Shapley value approximation method that trains an interpreter to get marginal Shapley values using a single forward pass.
  • VAEAC [17]: A generative model which obtains the conditional Shapley value by calculating the conditional expectation of the contribution function.
Since the true Shapley value of the real-world datasets are difficult to calculate, a quantitative metric is used to evaluate the performance of feature attributions.
Softmax information curve(SIC AUC) [48]. This metric evaluates the impact of attributed features on model output. Effective feature attribution should align closely with the model’s primary concerns [49]. The metric consists of two distinct indicators. The first indicator involves incrementally replacing the explained features with background features. Features with the highest attribution are introduced first, while those with the lowest attribution are added last, based on a predefined attribution threshold. Features with higher attribution values generally lead to more substantial improvements in model output performance. The model’s output curve can be plotted using the Softmax function, and the area under this curve can be quantified using the SIC AUC-ADD index. Conversely, the second indicator progressively removes features that contribute most to the output until all features are replaced by the background [50]. In this case, features with higher attribution values typically cause a more rapid decline in model performance, and the area under the resulting output curve is quantified using the SIC AUC-DEL index. A more accurate explanation yields a larger SIC AUC-ADD and a smaller SIC AUC-DEL index.
Hyperparameter selection FastSHAP, VAEAC, and EmSHAP are based on neural networks, and their hyperparameter are summarized in Table II. All MLPs in the table include two residual connections. All models are trained using the Adam optimizer with a learning rate of 0.001. For the MNIST dataset, the batch size for all three models is set as 256. For the TREC and ADTI datasets, the batch size is set as 64, and the batch size of ETT is set to be 100. For EmSHAP, the minimum and maximum thresholds for dynamic masking are set as $\zeta_{\min}=0.2$ and $\zeta_{max}=0.8$, respectively, and the dimension of context vector $\gamma$ is set as 32. It is worth noting that increasing the latent dimensions may improve the accuracy of explanations, but it will also significantly increase computational cost. To achieve a trade-off between accuracy and computational efficiency, and to ensure a fair comparison, the hidden layer size is set as 32 for VAEAC and EmSHAP, and 128 for FastSHAP.
TABLE II: Network structure of feature attribution methods on real-world datasets
Method Network Prototype Latent
dimensions
FastSHAP Surrogate MLP 128
Explainer MLP 128
VAEAC Full encoder MLP 32
Masked encoder MLP 32
Decoder MLP 32
EmSHAP Proposal network GRU $2|D|+|\bm{\gamma}|$
Energy network MLP 32
Table III summarizes the feature attribution performance and computation efficiency of different methods on four real-world datasets. All experiments are performed on a workstation equipped with an Intel Xeon Gold 6326 processor, 377G memory, and an RTX4080 graphics processing unit. The calculations are performed using Python 3.9 and PyTorch 2.4.1. The calculation time and memory usage for all methods are also compared and listed in Table III. From Table III it can be seen that EmSHAP yields the highest SIC AUC-ADD and the lowest SIC AUC-DEL in all cases, indicating that EmSHAP has the best feature attribution performance. For computation efficiency, Table III shows that methods that do not involve training stage are more efficient. However, for methods that involving both training and inference stages, it can be seen that EmSHAP has the highest efficiency as it requires the least amount of time.
TABLE III: Performance comparison of feature attribution for competitive methods on real-world datasets
Dataset Explainer SIC AUC Run time Memory usage
[…]
MNIST SamplingSHAP 0.893 0.269 /// 2.55s /// $\mathbf{1464.81}$MB
SVARM 0.257 0.125 /// 2.03s /// 1647.98MB
KernelSHAP 0.861 0.134 /// 1.84s /// 1714.42MB
FastSHAP 0.861 0.205 3302.66s $\mathbf{0.19}$s 2900.80MB 1958.56MB
VAEAC 0.714 0.447 5079.43s 2.71s 3373.30MB 1533.63MB
EmSHAP $\mathbf{0.965}$ $\mathbf{0.037}$ $\mathbf{2977.10}$s 1.63s $\mathbf{1823.91}$MB 1879.68MB
TREC SamplingSHAP 0.888 0.155 /// 1.62s /// 978.40MB
SVARM 0.846 0.187 /// 1.23s /// 980.22MB
KernelSHAP 0.875 0.135 /// 0.34s /// 993.72MB
FastSHAP 0.815 0.194 204.75s $\mathbf{0.14}$s 1168.51MB $\mathbf{890.36}$MB
VAEAC 0.725 0.316 646.19s 3.10s 1091.89MB 1071.34MB
EmSHAP $\mathbf{0.948}$ $\mathbf{0.103}$ $\mathbf{162.75}$s 0.19s $\mathbf{1085.62}$MB 994.82MB
ADTI SamplingSHAP 0.925 0.494 /// 0.20s /// 718.47MB
SVARM 0.927 0.518 /// $\mathbf{0.18s}$ /// 696.97MB
KernelSHAP 0.913 0.465 /// 1.43s /// 703.79MB
FastSHAP 0.914 0.535 295.87s 0.26s $\mathbf{717.17}$MB $\mathbf{696.64}$MB
VAEAC 0.855 0.596 804.22s 2.30s 1084.62MB 956.01MB
EmSHAP $\mathbf{0.933}$ $\mathbf{0.411}$ $\mathbf{270.90}$s 0.87s 966.58MB 946.61MB
ETT SamplingSHAP 0.893 0.526 /// 0.08s /// 761.11MB
SVARM 0.875 0.342 /// 0.05s /// 758.97MB
KernelSHAP 0.907 0.289 /// 7E-3s /// $\mathbf{534.54}$MB
FastSHAP 0.863 0.402 19.39s $\mathbf{1E-3}$s 707.49MB 586.20MB
VAEAC 0.902 0.359 46.91s 0.33s $\mathbf{530.85}$MB 568.14MB
EmSHAP $\mathbf{0.915}$ $\mathbf{0.243}$ $\mathbf{16.55}$s 4E-3s 554.11MB 540.46MB
VI-A Feature attribution for MNIST image classification
In this subsection, the MNIST dataset, which is widely used for evaluating XAI methods [14, 29, 11], is considered to test the performance of EmSHAP. The MNIST dataset provides a standardized benchmark for assessing algorithm performance in tasks such as classification and pattern recognition, it consists of 70,000 grayscale images of digits (0-9), each having a size of 28x28 pixels, with 60,000 images allocated for training and 10,000 for testing.
To classify the MNIST images, a simple CNN model is established, comprising two convolutional, two max-pooling layers and a fully connected output layer. The trained model achieved a classification accuracy of 97%, indicating a good performance. Given that CNN is a black-box model, it is difficult to understand the importance of different pixels. To perform this task, EmSHAP is employed to quantify the influence of input pixels on the classification results. It is trained and utilized to estimate the conditional probability distributions of the test data, which are further used to estimate the Shapley value.
In order to enhance computational efficiency, we use average pooling to reduce the 28×28 image to 14×14 image, and the number of attributing features is reduced to 196. Fig 4 shows the feature attribution results using EmSHAP and 5 competitive methods.
Figure 4: Feature attributions on MNIST dataset
As is shown in Fig 4, SVARM may face challenges in terms of interpretability when applied to image classification tasks, as it yields ambiguous results. Similarly, FastSHAP and VAEAC also produce ambiguous reconstructions. On the other hand, SamplingSHAP, KernelSHAP and EmSHAP all successfully attribute the digits of 3 and 8. A more delicate inspection, however, shows that SamplingSHAP and KernelSHAP yielded attribution that align with the classified digits, but tend to omit finer details. In contrast, EmSHAP exhibits finer details in feature attribution, especially for digit 3. The better performance is also verified by the higher SIC AUC-ADD value and lower SIC AUC-DEL value in Table III.
To test the performance of the proposed method under noisy data, Gaussian distributed noises with zero mean and varying degree of variances are added to the original images. The variances vary from 0.1 to 0.5, with an interval of 0.1. Feature attribution is then performed on these noisy images. Figs.5 and 6 present the attribution results produced by the six methods on the noisy images.
Figure 5: Feature attributions for noise-polluted image of digit 3. From top to bottom: noise variances of 0.1, 0.2, 0.3, 0.4, and 0.5.
Figure 6: Feature attributions for noise-polluted image of digit 8. From top to bottom: noise variances of 0.1, 0.2, 0.3, 0.4, and 0.5.
It can be seen from Figs.5 and 6 that when the variances of noise increase, the feature attribution performance of all methods deteriorates. Whilst the sharpest performance deterioration can be observed for SamplingSHAP, the performance for KernelSHAP and EmSHAP are relatively stable. This verifies the better robustness of EmSHAP in the case of noises.
VI-B Feature attribution for TREC question classification
The semantic classification of questions is a critical task in natural language processing, particularly in question-answering systems. Questions can be categorized into semantic classes based on their nature, such as inquiries about numbers, places, or objects. This classification aids in identifying potential answers. For instance, the question ”When did CNN begin broadcasting?” pertains to date and should be classified under the [Numeric] category.
In this subsection, the TREC question dataset, which includes six semantic classes, is used to evaluate the performance of EmSHAP. A CNN-based classifier (TextCNN) is used for classification [51], with a classification accuracy of 87.70%. For feature attribution, word-level attribution is performed to identify the trigger words that are most influential in determining the answer type.
Fig 7 presents a selection of questions randomly sampled from six distinct categories in the TREC test set. These questions are analyzed using the six methods to attribute word-level features to the trained TextCNN model. Compared to other methods, EmSHAP produces explanation results that better align with human language perception. The key advantages of EmSHAP over the other methods are summarized as follows: 1) Reduced attribution of generalized question words and question marks; 2) Increased attribution of key phrases; 3) Reduced attribution of connectives.
Figure 7: Feature attributions on TREC dataset
1) Reduced attribution of generalized question words and question marks: Certain key question words exhibit a strong correlation with question classification. For example, questions starting with ”Who” are classified as [Human] (question #4), those beginning with ”Where” as [Location] (question #5), and those starting with ”When” as [Numeric] (question #6). In these cases, the word-level feature attribution should reflect the prominence of these question words, meaning their Shapley values should be the highest. Conversely, the question word ”What” can initiate a wide range of question types (question #1–#3), making it weakly correlated with any specific category, and thus, less informative for question classification. Therefore, its Shapley value should be lower. Additionally, the presence of a question mark at the end of the sentence should not influence the classification, and as such, the Shapley value of the question mark should also be small.
As shown in Fig 7, for questions #1–#3, which begin with ”What,” EmSHAP reduces the contribution of ”What” and ”?” in feature attribution, focusing more on the contribution of the central phrase to question classification. This approach aligns with human cognitive processes. In contrast, other methods exhibit varying degrees of bias in attributing features to ”What” and ”?”.
2) Increased Attribution of Key Phrases: Compared to other methods, EmSHAP places greater emphasis on key phrases relevant to question classification. For instance, ”acronym NASA” corresponds to the category [Abbreviation], ”Phi Beta Kappa mean” corresponds to [Description], and ”popular toy in 1957” corresponds to [Entity].
3) Reduced Attribution of Connectives: Articles, prepositions, conjunctions, and similar words function to modify, connect, or supplement nouns or verbs, contributing to grammatical structures and the expression of complex meanings. Since these words do not stand alone in a sentence and are highly context-dependent, their contribution to question classification is relatively low in word-level feature attribution. Reducing the attribution of such words and prioritizing more essential short phrases is consistent with human semantic processing. As shown in Figure 7, EmSHAP reduces the attribution of words like ”the”, ”in” in [Entity] (question #3), ”in the” in [Human] (question #4), and ”are the” in [Location] (question #5).
VI-C Feature attribution for ADTI tabular classification
Classifying adult income using census data is a typical tabular classification task and has been widely used to test XAI methods. In this subsection, we use this dataset to test the performance of EmSHAP on tabular data. The census data collects information on 48,842 independent individuals, involving 12 input variables related to individual income and a discrete output variable $y$, where $y=0$ means that the individual income is less than or equal to 50k, and $y=1$ means that the individual income is greater than 50k. After the individual data with missing values in the data are removed, 32,562 samples remain. 80% of these samples are randomly selected as the training set, and the remaining samples are used as the test set. To classify adult income, a three-layer MLP model is established, and the hidden layer size is set to 64. The model is trained using the training set and is optimized using the Adam optimizer with a learning rate of 0.001. The classification accuracy of the model can reach 84.75%, indicating that the model has good classification performance.
To explain this predictive model, all six methods are used to quantify the contribution of input features. Fig 8 shows the mean absolute Shapley values of each feature.
Figure 8: Feature attributions for ADTI dataset
Among the six Shapley value estimation methods, SamplingSHAP, KernelSHAP, and EmSHAP identify age and capital gains as the most influential features in model predictions. This is in line with practical scenarios. Since the two variables are strongly correlated with income. For example, as people age, they typically accumulate more work experience and higher income. Capital gains are also positively correlated with income, as they generally stem from investment activities. In contrast, other factors like sex, marital status, and relationship tend not to be so related to income. On the other hand, SVARM also highlights relationship and sex as important. FastSHAP acknowledges that capital gains contribute most significantly, but the contributions of other variables are nearly identical. VAEAC suggests that marital status and relationship have a greater impact on adult income. Hence, in the case of ADTI dataset, the performance of feature attributions obtained from EmSHAP is also among the best.
Figure 9: Feature attributions on ADTI dataset for different income classes
To clearly illustrate the contribution of each feature to model classifications, we have categorized the data into two groups: income $\leq 50$K and income $>50$K. Fig. 9 presents the mean absolute Shapley values for each feature across the two income categories. In Fig. 9, bar charts with slashes and circles represent the absolute average Shapley values for the classification of individuals with income $\leq 50$K and $>50$K, respectively. It can be observed that, for the income $\leq 50$K group, age has the greatest contribution to the model’s classification, while for the income $>50$K group, capital gains hold the greatest influence. This difference can be explained. As individuals with income $\leq 50$K are typically younger or have less work experience, by accumulating age and work experience, their income will increase. Therefore, in the low-income group, age plays a more significant role in feature attribution. On the other hand, individuals with higher incomes tend to have stronger investment capabilities and more available funds for investment, which results in higher capital gains. Consequently, capital gains are a stronger attributor for individuals with income $>50$K.
VI-D Feature attributions for ETT dataset
In this subsection, the electricity transformer dataset (ETT), which has been widely used in long-term and short-term time series prediction tasks [52, 53], is considered. The time series data in ETT dataset is typically non-stationary and contains a reasonable portion of outliers. The dataset consists of 6 power load features and an output feature of ”oil temperature”, sampled at different temporal granularities. Here only the hourly data ETTh1 is used, which contains a total of 17420 samples.
In the ETT dataset, the task is to make predictions of oil temperature based on the 6 power load features. This will help the operators to make corrective operations before the oil temperature gets out of the normal interval. For the purpose of constructing predictive model, the Informer model [53] is applied, with the first 80% of the dataset used as training data and the last 20% as test data. The Informer model performs well in the test data, achieving a mean square error of less than 0.02. Despite the high accuracy, it is still desirable to know which input feature is more response for the predictions, which can be determined through estimation of Shapley values using EmSHAP and the other 5 competitive methods.
The feature attribution results can be referred to in Table III, which shows that EmSHAP outperforms competitive methods in terms of the SIC AUC metric. For a clear display, Fig. 10 presents the heatmap of estimated Shapley values using EmSHAP for the test set. It can be seen from Fig. 10 that the first four variables, HUFL, HULL, MULL, and MULL are more responsible for the model predictions, whilst the contributions of LUFL and LULL are negligible. This is in accordance with the feature attribution results of other competitive methods.
Figure 10: Heatmap of estimating Shapley values using EmSHAP in the ETTh1 prediction task
In order to further test the robustness of EmSHAP under adversarial attacks, a series of adversarial samples with varying attack amplitudes were generated using the Fast Gradient Sign Method (FGSM) [54] and added to the original test data. Based on the new test data, feature attribution is again performed using EmSHAP, FastSHAP and VAEAC, and the results are shown in Table IV. The reason for not comparing with other methods is that they do not involve a training stage. It can be seen from Table IV that the SIC AUC-ADD values of all three methods decrease after adding the adversarial samples. A similar result can be observed for the SIC AUC-DEL values, which increased for all the three methods, indicating the deterioration of feature attribution performance. A more delicate observation on the SIC AUC-ADD and SIC AUC-DEL values, however, shows that EmSHAP consistently outperforms the other two methods across all tested amplitudes.
TABLE IV: SIC AUC results for different attack amplitude on ETT dataset
Explainer SIC AUC Attack amplitude
0.02 0.04 0.06 0.08 0.10
[…]
VI-E Performance analysis of EmSHAP under different model structure
In this subsection, the performance of EmSHAP under different model structure is compared and analyzed.
VI-E1 Ablation study
An ablation study is performed to reflect the performance of EmSHAP on different standardized models, which involves the combinations of vanilla EBM, the GRU network, and the dynamic masking mechanism.
Table V presents the feature attribution performance and runtime under different cases. As is shown in Table V, although EmSHAP requires more time for training, it demonstrates superior performance in terms of SIC AUC metric. More specifically, the vanilla EBM yields the worst performance, by gradually adding the dynamic masking mechanism and GRU network, the performance increases. This clearly shows that the combined effect of these mechanisms contributes to the better performance of EmSHAP.
TABLE V: Results of ablation experiments
Dataset Explainer SIC AUC Run time Memory usage
[…]
MNIST Vanilla EBM 0.883 0.217 $\mathbf{1568.18}$s $\mathbf{1.02}$s $\mathbf{1354.19}$MB $\mathbf{1200.41}$MB
Vanilla EBM+Dynamic masking 0.900 0.173 1839.42s 1.12s 1439.98MB 1207.33MB
Vanilla EBM+GRU coupling 0.954 0.103 2840.98s 1.62s 1777.21MB 1730.01MB
EmSHAP $\mathbf{0.965}$ $\mathbf{0.037}$ 2977.10s 1.63s 1823.91MB 1879.68MB
TREC Vanilla EBM 0.808 0.213 $\mathbf{94.88}$s $\mathbf{0.14}$s $\mathbf{995.57MB}$ $\mathbf{923.68}$MB
Vanilla EBM+Dynamic masking 0.814 0.196 102.00s 0.15s 1055.01MB 989.62MB
Vanilla EBM+GRU coupling 0.922 0.142 156.33s 0.17s 1070.90MB 990.03MB
EmSHAP $\mathbf{0.948}$ $\mathbf{0.103}$ 162.75s 0.19s 1085.62MB 994.82MB
ADTI Vanilla EBM 0.863 0.544 $\mathbf{173.22}$s $\mathbf{0.81}$s $\mathbf{889.21}$MB $\mathbf{850.41}$MB
Vanilla EBM+Dynamic masking 0.872 0.469 184.36s 0.82s 950.21MB 892.53MB
Vanilla EBM+GRU coupling 0.927 0.420 253.74s 0.90s 963.37MB 930.10MB
EmSHAP $\mathbf{0.933}$ $\mathbf{0.411}$ 270.90s 0.87s 966.58MB 946.81MB
ETT Vanilla EBM 0.864 0.440 $\mathbf{14.19}$s $\mathbf{1E-3}$s $\mathbf{515.72}$MB $\mathbf{510.33}$MB
Vanilla EBM+Dynamic masking 0.873 0.339 14.77s 3E-3s 520.69MB 512.53MB
Vanilla EBM+GRU couping 0.903 0.292 15.99s 4E-3s 553.37MB 540.10MB
EmSHAP $\mathbf{0.915}$ $\mathbf{0.243}$ 16.55s 4E-3s 554.11MB 540.46MB
VI-E2 GRU network vs LSTM network
Table V shows that inclusion of GRU as the proposal network enhances the performance of EmSHAP. Since LSTM has a similar structure as GRU and might better capture long-term dependencies between variables, it makes sense to see how EmSHAP performs by replacing the GRU network with the LSTM network. For a fair comparison, the sizes of hidden layer for both LSTM and GRU are set as 32. Table VI summarizes the SIC AUC metric, runtime and memory using under the two cases. Table VI shows that using GRU as the proposal network yields more satisfactory results. Moreover, due to its simpler structure, GRU requires less runtime and memory usage. The fewer parameters and simpler structure of GRU reduce the risk of overfitting, which may explain why the SIC AUC results of GRU surpass those of LSTM.
TABLE VI: Proposal network in EmSHAP: LSTM vs GRU
Dataset Proposal network SIC AUC Run time Memory usage
[…]
MNIST LSTM 0.952 0.108 3406.11s 2.01s 2067.50MB 1999.63MB
GRU 0.965 0.037 2977.10s 1.63s 1879.91MB 1823.68MB
TREC LSTM 0.943 0.072 265.34s 0.22s 1201.12MB 1190.92MB
GRU 0.948 0.103 162.75s 0.19s 1085.62MB 994.82MB
ADTI LSTM 0.930 0.416 303.00s 0.91s 995.50MB 932.71MB
GRU 0.933 0.411 270.90s 0.87s 966.58MB 946.81MB
ETT LSTM 0.906 0.292 17.39s 6E-3s 565.50MB 552.71MB
GRU 0.915 0.243 16.55s 4E-3s 554.11MB 540.46MB
VI-E3 Dynamic masking vs fixed-rate masking
As an important component of EmSHAP, the dynamic masking mechanism enabling the model performance. To better show the functionality of dynamic masking, the performance of EmSHAP under dynamic masking and fix-rate masking is compared and the results are shown in Table VII.
TABLE VII: SIC AUC results for different masking rate on four datasets
Dataset SIC AUC Masking rate
0.5 0.6 0.7 0.8 0.9 Dynamic
masking
[…]
As shown in Fig. VII, the dynamic masking achieves better performance than fixed-rate masking for all cases. In addition, using the fixed-rate masking requires manually tuning the optimal masking rate, which is undesired in practice. In contrast, the dynamic masking mechanism allows the model to learn masking rate automatically, which leads to the better performance [55].
VII Conclusion
This paper proposes an energy-based model for Shapley value estimation (EmSHAP) for feature attribution in deep learning-based model. In EmSHAP, an energy-based model (EBM) is used to estimate the Shapley contribution function under different feature combinations. The EBM consists of a Resnet structure for approximating the energy function and a GRU network for estimating the partition distribution. By mapping the input features into a latent space using the GRU network, the influence of feature ordering to calculation of contribution function can be eliminated. In order to explore a wider range of feature combinations, a dynamic masking mechanism is introduced during training, which empowers EmSHAP with better estimation accuracy and generalization capability. In order to evaluate the performance of the proposed method, detailed theoretical analysis on error bound and extensive application studies are conducted, which shows that EmSHAP has tighter error bound and exhibits higher accuracy compared with state-of-the-art approaches. It should be noted that despite the theoretically universal approximation capability, EmSHAP may not be suitable to handle the small-sample problem. In addition, although EmSHAP exhibits better performance under conditions like noises, outliers, and adversarial perturbations, it still lacks an inherent mechanism to deal with these cases. Future work will focus on addressing these limitations to enhance the robustness and scalability.
Appendix
In this appendix, the error bounds of KernerlSHAP and VAEAC are analyzed.
VII-A Error Bound of KernelSHAP
In KernelSHAP, the calculation of Shapley value is achieved using an additive model, with the contribution function $v(S)$ of the variable set $S$ approximated as follows.
$\small\hat{v}(S)=\beta_{0}+\sum_{i\in S}\beta_{i},$ (A1)
where $\beta_{0}$ is the expectation of the deep learning model outputs, $\beta_{i}$ is the Shapley value of the $i$-th variable. KernelSHAP uses a linear regression model to obtain the Shapley value of each variable, which is estimated by solving the weighted least squares problem as follows.
$\small\mathop{\min}_{\beta_{0},...,\beta_{|D|}}\sum_{S\in D}\psi(S)\left(\hat{v}(S)-v(S)\right)^{2},$ (A2)
Here $\psi(S)\in[0,1]$ is called the Shapley kernel. To solve the above optimization problem, Covert et al. [12] introduced a binary vector $\mathbf{b}\in\mathbb{R}^{|D|+1}$ to determine whether the $i$-th feature is included in the feature subset $S$. Specifically, $b_{i}=1$ indicates the inclusion of $i$-th feature in $S$, while $b_{i}=0$ represents its exclusion. Thus, Eq.(A2) can be rewritten as follows.
$\small\mathop{\min}_{\bm{\beta}}\sum_{S\in D}\psi(S)\left(\mathbf{b}^{T}\bm{\beta}-v(S)\right)^{2}.$ (A3)
Here, $\bm{\beta}=[\beta_{0},...,\beta_{|D|}]^{T}\in\mathbb{R}^{|D|+1}$. The matrix form of Eq.(A3) is as follows.
$\small\mathop{\min}_{\bm{\beta}}\frac{1}{K}(\mathbf{B}^{T}\bm{\beta}-\mathbf{v})^{T}\bm{\psi}(\mathbf{B}^{T}\bm{\beta}-\mathbf{v}),$ (A4)
where $K$ is the number of sampling times, $\mathbf{B}\in\mathbb{R}^{(|D|+1)\times K}$ is a binary matrix indicating which variables are included in $S$, $\bm{\psi}\in\mathbb{R}^{K\times K}$ is a diagonal matrix with Shapley kernel in its diagonal, $\mathbf{v}\in\mathbb{R}^{K\times 1}$ is the contribution function vector.
The optimal estimation $\hat{\bm{\beta}}$ of Eq.(A4) can be obtained by taking the first-order derivative of $\bm{\beta}$ in Eq.(A4) and setting it to zero.
$\small\frac{1}{K}\left(\mathbf{B}\bm{\psi}\mathbf{B}^{T}{\bm{\beta}}-\mathbf{B}^{T}\bm{\psi}\mathbf{v}\right)=0.$ (A5)
For more details, interested readers can refer to Refs. [11, 12].
Next we will analyze the error bound of KernelSHAP. Let $\hat{\bm{\Sigma}}=\frac{1}{K}\mathbf{B}\bm{\psi}\mathbf{B}^{T}$ and $\hat{\bm{\Gamma}}=\frac{1}{K}\mathbf{B}^{T}\bm{\psi}\mathbf{v}$, $\mathbb{E}[\hat{\bm{\Sigma}}]=\bm{\Sigma}^{*}$ and $\mathbb{E}[\hat{\bm{\Gamma}}]=\bm{\Gamma}^{*}$. Assuming that $\hat{\bm{\Sigma}}$ and $\bm{\Sigma}^{*}$ are invertible, we have $\hat{\bm{\beta}}=\hat{\bm{\Sigma}}^{-1}\hat{\bm{\Gamma}}$, $\bm{\beta}^{*}=\mathbb{E}[\hat{\bm{\Sigma}}^{-1}\hat{\bm{\Gamma}}]$. The error bound of KernelSHAP can be described as the mean absolute deviation between the estimated contribution function $\hat{v}(S)$ and the true contribution function $v(S)$, which is related to $\hat{\bm{\beta}}$ and $\bm{\beta}^{*}$ as follows.
$\small\mathbb{E}\left[|\hat{v}(S)-v(S)|\right]\leq\mathbb{E}\left[||\mathbf{b}^{T}||\cdot||\hat{\bm{\beta}}-\bm{\beta}^{*}||\right]=(|S|+1)\mathbb{E}\left[||\hat{\bm{\beta}}-\bm{\beta}^{*}||\right].$ (A6)
The decomposition of $\mathbb{E}\left[||\hat{\bm{\beta}}-\bm{\beta}^{*}||\right]$ is given by,
$\quad\mathbb{E}\left[||\hat{\bm{\beta}}-\bm{\beta}^{*}||\right]=\mathbb{E}[||\hat{\bm{\Sigma}}^{-1}\hat{\bm{\Gamma}}-\mathbb{E}[\hat{\bm{\Sigma}}^{-1}\hat{\bm{\Gamma}}]||]$ (A7)
$=\mathbb{E}[||\hat{\bm{\Sigma}}^{-1}\hat{\bm{\Gamma}}-\mathbb{E}[\hat{\bm{\Sigma}}^{-1}]\mathbb{E}[\hat{\bm{\Gamma}}]-\text{Cov}(\hat{\bm{\Sigma}}^{-1},\hat{\bm{\Gamma}})||]$
$=\mathbb{E}[||\hat{\bm{\Sigma}}^{-1}\hat{\bm{\Gamma}}-\bm{\Sigma}^{*^{-1}}\bm{\Gamma}^{*}-\text{Cov}(\hat{\bm{\Sigma}}^{-1},\hat{\bm{\Gamma}})||],$
where |||||$\cdot$||| | ⋅ | | is the $l_{2}$ norm and $\text{Cov}(\cdot,\cdot)$ is the covariance. Next we introduce three corollaries and show that $||\text{Cov}(\hat{\bm{\Sigma}}^{-1},\hat{\bm{\Gamma}})||\leq\epsilon_{3}^{2}$ holds in some specific cases, with $\epsilon_{3}$ being a small positive number.
Corollary A1.
(Concentration of $\hat{\bm{\Sigma}}$) For any $\epsilon_{3}>0$, the following holds,
$\small\mathbb{P}(||\hat{\bm{\Sigma}}-\bm{\Sigma}^{*}||\geq\epsilon_{3})\leq 2(|D|+1)^{2}\exp(-2K\epsilon_{3}^{2}).$ (A8)
Proof. Noting that the elements in $\mathbf{B}$ are 0 or 1 and the elements in $\bm{\psi}$ is bounded within $[0,1]$. It can be seen that the elements in $\hat{\bm{\Sigma}}$ and $\bm{\Sigma}^{*}$ are also bounded within $[0,1]$. According to Hoeffdings inequality, each element in $\hat{\bm{\Sigma}}$ is bounded with their expectation values, that is, for any $\epsilon_{3}>0$
$\small\begin{cases}\mathbb{P}(|\frac{1}{K}\sum_{k=1}^{K}(\hat{\bm{\Sigma}}_{ii}-\mathbb{E}[\hat{\bm{\Sigma}}_{ii}])|\geq\epsilon_{3})\leq 2\exp(-2K\epsilon_{3}^{2})\\ \mathbb{P}(|\frac{1}{K}\sum_{k=1}^{K}(\hat{\bm{\Sigma}}_{ij}-\mathbb{E}[\hat{\bm{\Sigma}}_{ij}])|\geq\epsilon_{3})\leq 2\exp(-2K\epsilon_{3}^{2}),\end{cases}$ (A9)
By using Booles inequality [56], Eq.(A8) can be obtained.
The concentration of $\hat{\bm{\Gamma}}$ and $\hat{\bm{\Sigma}}^{-1}$ can be obtained in the same.
Corollary A2.
(Concentration of $\hat{\bm{\Gamma}}$)For any $\epsilon_{3}>0$, the following inequality holds.
$\small\mathbb{P}(||\hat{\bm{\Gamma}}-\bm{\Gamma}^{*}||\geq\epsilon_{3})\leq 2(|D|+1)\exp(-2K\epsilon_{3}^{2}).$ (A10)
Proof. According to Assumption 4, the contribution value of any subset $S$ is bounded within $[0,1]$, that is $v(S)\in[0,1]$. Then $\mathbf{v}$ is also bounded within $[0,1]$ and $\bm{\Gamma}^{*}\in[0,1]$. By using Hoeffdings inequality, each element in $\hat{\bm{\Gamma}}$ is bounded with their expectation values.
$\small\mathbb{P}(|\frac{1}{K}\sum_{k=1}^{K}(\hat{\bm{\Gamma}}_{i}-\mathbb{E}[\hat{\bm{\Gamma}}_{i}])|\geq\epsilon_{3})\leq 2\exp(-2K\epsilon_{3}^{2}).$ (A11)
By using Booles inequality [56], the concentration of $\hat{\bm{\Gamma}}$ in Eq.(A10) can be obtained.
Corollary A3.
(Concentration of $\hat{\bm{\Sigma}}^{-1}$) For any $\epsilon_{3}>0$, we have,
$\small\mathbb{P}(||\hat{\bm{\Sigma}}^{-1}-{\bm{\Sigma}^{*}}^{-1}||\geq\epsilon_{3})\leq 2(|D|+1)^{2}\exp(-2K\epsilon_{3}^{2}).$ (A12)
Corollary A4.
For a fixed input dimension $|D|$ and any small positive number $\epsilon_{3}$, there exists a large $K$ that
$\small\mathbb{P}\left(||\mathrm{Cov}(\hat{\bm{\Sigma}}^{-1},\hat{\bm{\Gamma}})||\geq\epsilon^{2}_{3}\right)\leq 2((|D|+1)^{2}+(|D|+1))\exp(-2K\epsilon_{3}^{2}),$ (A13)
where $\epsilon_{3}\in\left[\right.\sqrt{\frac{1}{2K}\log(2((|D|+1)^{2}+(|D|+1)))},+\infty\left.\right)$.
Proof. $\left|\left|\text{Cov}(\hat{\bm{\Sigma}}^{-1},\hat{\bm{\Gamma}})\right|\right|$ can be decomposed as follows.
$\small\quad\left|\left|\text{Cov}(\hat{\bm{\Sigma}}^{-1},\hat{\bm{\Gamma}})\right|\right|=\left|\left|\mathbb{E}\left[\left(\hat{\bm{\Sigma}}^{-1}-\mathbb{E}[\hat{\bm{\Sigma}}^{-1}]\right)\left(\hat{\bm{\Gamma}}-\mathbb{E}[\hat{\bm{\Gamma}}]\right)\right]\right|\right|.$ (A14)
According to Corollaries A2, A3, and Booles inequality, Eq.(A13) can be obtained. Since the probability $\mathbb{P}\left(||\text{Cov}(\hat{\bm{\Sigma}}^{-1},\hat{\bm{\Gamma}})||\geq\epsilon_{3}\right)\in[0,1]$, the range of $\epsilon_{3}$ is $\left[\right.\sqrt{\frac{1}{2K}\log(2((|D|+1)^{2}+(|D|+1)))},+\infty\left.\right)$.
It should be noted that for a small positive $\epsilon_{3}$, when $K$ is large, the probability $1-2((|D|+1)^{2}+(|D|+1))\exp(-2K\epsilon_{3}^{2})$ is close to 1, indicating that $\left|\left|\text{Cov}(\hat{\bm{\Sigma}}^{-1},\hat{\bm{\Gamma}})\right|\right|\leq\epsilon_{3}^{2}$ holds true under almost any conditions.
Thus, Eq.(A7) can be rewritten as follows,
$\quad\mathbb{E}\left[||\hat{\bm{\beta}}-\bm{\beta}^{*}||\right]$ (A15)
$\leq\mathbb{E}[||\hat{\bm{\Sigma}}^{-1}\hat{\bm{\Gamma}}-\bm{\Sigma}^{*^{-1}}\bm{\Gamma}^{*}||]+\mathbb{E}[||-\text{Cov}(\hat{\bm{\Sigma}}^{-1},\hat{\bm{\Gamma}})||]$
$\leq\mathbb{E}[||\hat{\bm{\Sigma}}^{-1}(\hat{\bm{\Gamma}}-\bm{\Gamma}^{*})||]+\mathbb{E}[||(\hat{\bm{\Sigma}}^{-1}-\bm{\Sigma}^{*^{-1}})\bm{\Gamma}^{*}||]+\epsilon_{3}^{2}$
$\leq\mathbb{E}[||\hat{\bm{\Sigma}}^{-1}||\cdot||\hat{\bm{\Gamma}}-\bm{\Gamma}^{*}||]+\mathbb{E}[||\hat{\bm{\Sigma}}^{-1}-\bm{\Sigma}^{*^{-1}}||\cdot||\bm{\Gamma}^{*}||]+\epsilon_{3}^{2}$
$=\mathbb{E}[||\hat{\bm{\Sigma}}^{-1}(\hat{\bm{\Gamma}}-\bm{\Gamma}^{*})||]+\mathbb{E}[||\bm{\Sigma}^{*^{-1}}(\hat{\bm{\Sigma}}-\bm{\Sigma}^{*})\hat{\bm{\Sigma}}^{-1}||\cdot||\bm{\Gamma}^{*}||]+\epsilon_{3}^{2}$
$\leq\mathbb{E}[||\hat{\bm{\Sigma}}^{-1}||\cdot||\hat{\bm{\Gamma}}-\bm{\Gamma}^{*}||]+\mathbb{E}[||\bm{\Sigma}^{*^{-1}}||\cdot||\hat{\bm{\Sigma}}-\bm{\Sigma}^{*}||\cdot||\hat{\bm{\Sigma}}^{-1}||\cdot||\bm{\Gamma}^{*}||]+\epsilon_{3}^{2}.$
By calculating the upper bounds of the decomposed elements in Eq.(A15) separately and merging them, the upper bound of $\mathbb{E}\left[|\hat{v}(S)-v(S)|\right]$ can be obtained. Notice that Ref.[57] gives the theoretical analysis of LIME, we follow their contribution and derive the error bound of KernelSHAP. Here, the explicit expressions of $\bm{\Sigma}^{*}$ and its inverse form $\bm{\Sigma}^{*}$ are first taken into consideration. According to Ref. [12], $\bm{\Sigma}^{*}$ has the following form,
$\small\bm{\Sigma}^{*}=\begin{bmatrix}\frac{1}{2}&\alpha&\cdots&\alpha\\ \alpha&\frac{1}{2}&\cdots&\alpha\\ \vdots&\vdots&\ddots&\vdots\\ \alpha&\alpha&\cdots&\frac{1}{2},\end{bmatrix}$ (A16)
that is, the diagonal of $\bm{\Sigma}^{*}$ is $\frac{1}{2}$, and the off-diagonal entries $\alpha$ are computed as follows.
$\small\alpha=\frac{1}{|D|(|D|+1)}\frac{\sum_{i=2}^{|D|}\frac{i-1}{|D|+1-i}}{\sum_{i=1}^{|D|}\frac{1}{i(|D|+1-i)}}.$ (A17)
Lemma A1.
Assuming $\bm{\Sigma}^{*}$ is invertible, the inverse of $\bm{\Sigma}^{*}$, $\bm{\Sigma}^{*^{-1}}$, has the following expression,
$\small\bm{\Sigma}^{*^{-1}}=\begin{bmatrix}\frac{1}{\frac{1}{2}-\alpha}&\frac{\alpha}{(\alpha-\frac{1}{2})(K\alpha-\alpha+\frac{1}{2})}&\cdots&\frac{\alpha}{(\alpha-\frac{1}{2})(K\alpha-\alpha+\frac{1}{2})}\\ \frac{\alpha}{(\alpha-\frac{1}{2})(K\alpha-\alpha+\frac{1}{2})}&\frac{1}{\frac{1}{2}-\alpha}&\cdots&\frac{\alpha}{(\alpha-\frac{1}{2})(K\alpha-\alpha+\frac{1}{2})}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\alpha}{(\alpha-\frac{1}{2})(K\alpha-\alpha+\frac{1}{2})}&\frac{\alpha}{(\alpha-\frac{1}{2})(K\alpha-\alpha+\frac{1}{2})}&\cdots&\frac{1}{\frac{1}{2}-\alpha}\end{bmatrix}.$ (A18)
Proof. According to Eq.(A16), $\bm{\Sigma}^{*}$ can be decomposed as follows,
$\small\bm{\Sigma}^{*}=(\frac{1}{2}-\alpha)\mathbf{I}+\alpha\mathbf{J},$ (A19)
where $\mathbf{I}\in\mathbb{R}^{(|D|+1)\times(|D|+1)}$ is an identity matrix and $\mathbf{J}\in\mathbb{R}^{(|D|+1)\times(|D|+1)}$ is a matrix with all ones. The inverse of $\bm{\Sigma}^{*}$ can be denoted as $\bm{\Sigma}^{*{-1}}=\kappa\mathbf{I}+\omega\mathbf{J}$.
$\small\left(\left(\frac{1}{2}-\alpha\right)\mathbf{I}+\alpha\mathbf{J}\right)\left(\kappa\mathbf{I}+\omega\mathbf{J}\right)=\mathbf{I}.$ (A20)
By sorting out the above formula,
$\small\left(\frac{1}{2}-\alpha\right)\kappa\mathbf{I}+\left[\left(\frac{1}{2}-\alpha+K\alpha\right)\omega+\kappa\alpha\right]\mathbf{J}=\mathbf{I}.$ (A21)
Thus, the following holds
[…]
By solving Eq.(A22), we have,
$\small\kappa=\frac{1}{\frac{1}{2}-\alpha},\quad\omega=\frac{\alpha}{(\alpha-\frac{1}{2})(K\alpha-\alpha+\frac{1}{2})},$ (A23)
and
$\small\bm{\Sigma}^{*^{-1}}=\frac{1}{\frac{1}{2}-\alpha}\mathbf{I}+\frac{\alpha}{(\alpha-\frac{1}{2})(K\alpha-\alpha+\frac{1}{2})}\mathbf{J}.$ (A24)
Proposition A1.
(The upper bound of $||\bm{\Sigma}^{*{-1}}||$) The $l_{2}$ norm of $\bm{\Sigma}^{*^{-1}}$ is upper bounded by,
$\small||\bm{\Sigma}^{*^{-1}}||\leq\sqrt{\frac{|D|+1}{(\frac{1}{2}-\alpha)^{2}}+\frac{(|D|+1)|D|\alpha^{2}}{((\alpha-\frac{1}{2})(K\alpha-\alpha+\frac{1}{2}))^{2}}}.$ (A25)
Proof. The square $l_{2}$ norm of $\bm{\Sigma}^{*^{-1}}$ is upper bounded by its Frobenius norm, that is,
$\small||\bm{\Sigma}^{*^{-1}}||^{2}\leq||\bm{\Sigma}^{*^{-1}}||^{2}_{F}=\frac{|D|+1}{(\frac{1}{2}-\alpha)^{2}}+\frac{(|D|+1)|D|\alpha^{2}}{((\alpha-\frac{1}{2})(K\alpha-\alpha+\frac{1}{2}))^{2}}.$ (A26)
By taking root square on both sides of Eq.(A26), we can get Eq.(A25).
The upper bound of $||\bm{\Gamma}^{*}||$ can be obtained in a same way.
Proposition A2.
(Upper bound of $||\bm{\Gamma}^{*}||$) The $l_{2}$ norm of $\bm{\Gamma}^{*}$ is upper bounded by,
$\small||\bm{\Gamma}^{*}||\leq\sqrt{|D|+1}.$ (A27)
Proof. The square $l_{2}$ norm of $\bm{\Gamma}^{*}$ is as follows.
$\small||\bm{\Gamma}^{*}||^{2}=||\mathbf{B}\bm{\psi}\mathbf{v}||^{2}.$ (A28)
Since each element in $\bm{\Gamma}^{*}$ is bounded within $[0,1]$, the square $l_{2}$ norm of $\bm{\Gamma}^{*}$ is smaller then its dimension, namely, $||\bm{\Gamma}^{*}||^{2}\leq|D|+1$.
Proposition A3.
(Upper bound of $\hat{\bm{\Sigma}}^{-1}$) For all $\epsilon_{3}\in(0,\frac{1}{||\bm{\Sigma}||_{F}^{*^{-1}}})$, with the probability greater than $1-2(|D|+1)^{2}\exp(-2K\epsilon_{3}^{2})$, there exist a small positive number $\epsilon_{3}^{\prime}$ that
$\small||\hat{\bm{\Sigma}}^{-1}||\leq\frac{||\bm{\Sigma}^{*^{-1}}||_{F}}{1-\epsilon_{3}||\bm{\Sigma}^{*^{-1}}||_{F}}=||\bm{\Sigma}^{*^{-1}}||_{F}+\epsilon_{3}^{\prime},$ (A29)
where $\epsilon_{3}^{\prime}=\frac{\epsilon_{3}||\bm{\Sigma}^{*^{-1}}||_{F}^{2}}{1-\epsilon_{3}||\bm{\Sigma}^{*^{-1}}||_{F}}$.
Proof. According to Proposition A1, we have $\lambda_{\max}(\bm{\Sigma}^{*^{-1}})\leq||\bm{\Sigma}^{*^{-1}}||_{F}$ and
$\small\lambda_{\min}(\bm{\Sigma}^{*})\geq\frac{1}{||\bm{\Sigma}^{*^{-1}}||_{F}}.$ (A30)
According to Corollary  A1, $\hat{\bm{\Sigma}}$ is concentrated on $\bm{\Sigma}^{*}$. By using Weyl inequality [58],
$\small|\lambda_{\min}(\hat{\bm{\Sigma}})-\lambda_{\min}(\bm{\Sigma}^{*})|\leq||\hat{\bm{\Sigma}}-\bm{\Sigma}^{*}||\leq\epsilon_{3},$ (A31)
so that,
$\small\lambda_{\min}(\hat{\bm{\Sigma}})\geq\lambda_{\min}(\bm{\Sigma}^{*})-\epsilon_{3}\geq\frac{1}{||\bm{\Sigma}^{*^{-1}}||_{F}}-\epsilon_{3},$ (A32)
which leads to
$\small||\hat{\bm{\Sigma}}^{-1}||\leq\frac{||\bm{\Sigma}^{*^{-1}}||_{F}}{1-\epsilon_{3}||\bm{\Sigma}^{*^{-1}}||_{F}}.$ (A33)
The right hand side of Eq.(A33) can bo obtained as follows.
$\quad\frac{||\bm{\Sigma}^{*^{-1}}||_{F}}{1-\epsilon_{3}||\bm{\Sigma}^{*^{-1}}||_{F}}=\frac{||\bm{\Sigma}^{*^{-1}}||_{F}-\epsilon_{3}||\bm{\Sigma}^{*^{-1}}||_{F}^{2}+\epsilon_{3}||\bm{\Sigma}^{*^{-1}}||_{F}^{2}}{1-\epsilon_{3}||\bm{\Sigma}^{*^{-1}}||_{F}}$ (A34)
$=||\bm{\Sigma}^{*^{-1}}||_{F}+\frac{\epsilon_{3}||\bm{\Sigma}^{*^{-1}}||_{F}^{2}}{1-\epsilon_{3}||\bm{\Sigma}^{*^{-1}}||_{F}}.$
By defining $\epsilon_{3}^{\prime}=\frac{\epsilon_{3}||\bm{\Sigma}^{*^{-1}}||_{F}^{2}}{1-\epsilon_{3}||\bm{\Sigma}^{*^{-1}}||_{F}}$, we can get the upper bound of $||\hat{\bm{\Sigma}}^{-1}||\leq|\hat{\bm{\Sigma}}^{-1}||_{F}+\epsilon_{3}^{\prime}$, with $\epsilon_{3}^{\prime}$ being a small number.
Proof. (Proof of Theorem 2) For fixed feature dimension $|D|$, a small positive $\epsilon_{3}$, and a large sampling number $K$, $\mathbb{E}\left[|\hat{v}(S)-v(S)|\right]$ is upper bounded with the probability greater than $1-2((|D|+1)^{2}+(|D|+1))\exp(-2K\epsilon_{3}^{2})$ as follows.
$\quad\mathbb{E}\left[|\hat{v}(S)-v(S)|\right]\leq\mathbb{E}\left[||\mathbf{b}^{T}||\cdot||\hat{\bm{\beta}}-\bm{\beta}^{*}||\right]=(|S|+1)\mathbb{E}\left[||\hat{\bm{\beta}}-\bm{\beta}^{*}||\right]$ (A35)
$\leq\{\mathbb{E}\left[||\hat{\bm{\Sigma}}^{-1}||\cdot||\hat{\bm{\Gamma}}-\bm{\Gamma}^{*}||\right]+\mathbb{E}\left[||\bm{\Sigma}^{*^{-1}}||\cdot||\hat{\bm{\Sigma}}-\bm{\Sigma}^{*}||\cdot||\hat{\bm{\Sigma}}^{-1}||\cdot||\bm{\Gamma}^{*}||\right]$
$\quad+\epsilon_{3}^{2}\}(|S|+1)$
$\leq\{||\bm{\Sigma}^{*^{-1}}||_{F}\cdot\int_{0}^{+\infty}2(|D|+1)\exp(-2K\epsilon_{3}^{2})d\epsilon_{3}+\epsilon_{3}^{2}$
[…]
$\cdot(||\bm{\Sigma}^{*^{-1}}||_{F}+\epsilon_{3}^{\prime})\cdot\sqrt{|D|+1}\}(|S|+1)$
$=2(|D|+1)(|S|+1)\frac{\sqrt{\pi}}{\sqrt{2K}}||\bm{\Sigma}^{*^{-1}}||_{F}$
$+2(|D|+1)^{\frac{5}{2}}(|S|+1)\frac{\sqrt{\pi}}{\sqrt{2K}}||\bm{\Sigma}^{*^{-1}}||_{F}(||\bm{\Sigma}^{*^{-1}}||_{F}+\epsilon_{3}^{\prime})+\epsilon_{3}^{2}.$
It can be seen from Eq.(A35) that as the data dimension $|D|$ increases, the error bound of KernelSHAP will increase. On the other hand, by increasing the number of sampling $K$ the error bound will get small.
VII-B Error Bound of VAEAC
Variational autoencoder (VAE) is a typical generative model, which uses a latent variable model to generate a probabilistic representation of the unknown data distribution $p(\mathbf{x})$. The basic idea of the variational autoencoder is to generate the observed variable $\mathbf{x}$ through a Gaussian distributed latent variable $\mathbf{z}$. Specifically, VAE models the joint distribution of $\mathbf{x}$ and $\mathbf{z}$ as follows.
$\small p(\mathbf{x},\mathbf{z})=p(\mathbf{z})p(\mathbf{x}|\mathbf{z}),$ (B1)
Here the conditional distribution $p(\mathbf{x}|\mathbf{z})$ is called an encoder, $\mathbf{z}$ is the latent variable. By marginalizing the latent variable, the data distribution can be approximated.
[…]
This latent variable model helps to get $p(\mathbf{x})$ easily. However, the calculation of the posterior distribution $p(\mathbf{z}|\mathbf{x})$ is always intractable. To address this challenge, VAE constructs a distribution $q(\mathbf{z}|\mathbf{x})$ to approximate $p(\mathbf{z}|\mathbf{x})$. The data log-likelihood of VAE can be expressed as follows.
[…]
$=\mathbb{E}_{q(\mathbf{z}|\mathbf{x})}\log p(\mathbf{x})$
$=\mathbb{E}_{q(\mathbf{z}|\mathbf{x})}\log\frac{p(\mathbf{x},\mathbf{z})}{p(\mathbf{z}|\mathbf{x})}$
$=\mathbb{E}_{q(\mathbf{z}|\mathbf{x})}\log\frac{p(\mathbf{x},\mathbf{z})q(\mathbf{z}|\mathbf{x})}{p(\mathbf{z}|\mathbf{x})q(\mathbf{z}|\mathbf{x})}$
$=\mathbb{E}_{q(\mathbf{z}|\mathbf{x})}\log\frac{p(\mathbf{x},\mathbf{z})}{q(\mathbf{z}|\mathbf{x})}+\mathbb{E}_{q(\mathbf{z}|\mathbf{x})}\log\frac{q(\mathbf{z}|\mathbf{x})}{p(\mathbf{z}|\mathbf{x})}$
$=\mathbb{E}_{q(\mathbf{z}|\mathbf{x})}\log\frac{p(\mathbf{x},\mathbf{z})}{q(\mathbf{z}|\mathbf{x})}+D_{KL}\left(q(\mathbf{z}|\mathbf{x})||p(\mathbf{z}|\mathbf{x})\right),$
where $\mathbb{E}_{q(\mathbf{z}|\mathbf{x})}\log\frac{p(\mathbf{x},\mathbf{z})}{q(\mathbf{z}|\mathbf{x})}$ is called evidence lower bound (ELBO) and $D_{KL}(\cdot||\cdot)$ is Kullback-Leibler (KL) divergence. ELBO is the objective function that needs to be optimized, by maximizing ELBO via gradient descent, the parameters of VAE can be obtained.
VAEAC, a Shapley value approximation approach based on VAE, is introduced by Ref. [17]. In VAEAC, the contribution function of Shapley value is sampled from the conditional probability $p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ generated by VAE. The log-likelihood introduced in Ref. [17] is as follows.
$\quad\log p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ (B4)
$=\mathbb{E}_{q(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}\left[\log\frac{p(\mathbf{x}_{\bar{S}},\mathbf{z}|\mathbf{x}_{S})}{q(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}\right]+\mathbb{E}_{q(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}\left[\log\frac{q(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}{p(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}\right]$
$=\mathbb{E}_{q(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}\left[\log\frac{p(\mathbf{x}_{\bar{S}},\mathbf{z}|\mathbf{x}_{S})}{q(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}\right]+D_{KL}(q(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})||p(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S}))$
$\geq\mathbb{E}_{q(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}\left[\log\frac{p(\mathbf{x}_{\bar{S}},\mathbf{z}|\mathbf{x}_{S})}{q(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}\right].$
VAEAC trains the parameters by maximizing the last term in Eq.(B4), and then samples from the generated conditional distribution $p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ to calculate the contribution function $\hat{v}(S)$.
According to Ref. [18], the inference error of VAE can be decomposed into two components, the approximation error and the amortization error. Since the analysis of the model error bound in this paper assumes the model can achieve theoretical optimal performance after a large number of iterative training, the approximation error of VAEAC, i.e. $\log p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})-p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$, can be controlled and upper bounded by a small positive number. Thus, similar to Assumption 1, the probability density difference between $p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ and $p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ is bounded as follows.
$\small\left|\frac{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-1\right|\leq\epsilon_{1},$ (B5)
Another non-negligible error of VAEAC is the amortization error, i.e. $\log p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})-\log p_{\theta}(\mathbf{x}_{\bar{S}|\mathbf{x}_{S}})$, which arises because VAE uses a shared encoder to approximate the posterior distribution for all data points, rather than optimizing a distinct posterior distribution for each individual sample. This amortization error also leads to the following difference between the probability densities.
$\small\left|\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-1\right|\leq\delta.$ (B6)
$\mathbb{E}[|\hat{v}(S)-v(S)|]$ of VAEAC can be derived as follows.
$\quad\mathbb{E}[|\hat{v}(S)-v(S)|]$ (B7)
$=\mathbb{E}\left[\left|\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|\right]$
$=\mathbb{E}\left[\left|\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right.\right.$
$\quad\left.\left.+\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right.\right.$
$\quad\left.\left.+\mathbb{E}_{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|\right]$
$\leq\mathbb{E}\left[\left|\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|\right]$
$\quad+\mathbb{E}\left[\left|\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right.\right.$
$\quad\left.\left.+\mathbb{E}_{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|\right].$
The first term in the last inequality of Eq.(B7) is the statistical error, the second and the third terms are related to the amortization error and approximation error of VAE, respectively. The upper bound of the statistical error is as follows.
$\small\mathbb{E}\left[\left|\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|\right]\leq\frac{\sqrt{\pi}}{\sqrt{2K}}.$ (B8)
The amortization error is bounded as follows.
$\quad\mathbb{E}\left[\left|\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|\right]$ (B9)
$=\mathbb{E}\left[\left|\int f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}-\int f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}\right|\right]$
$=\mathbb{E}\left[\left|\int f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}\right.\right.$
$\left.\left.\quad-\int f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}\right|\right]$
$=\mathbb{E}\left[\left|\mathbb{E}_{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\left(f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\right)\right]\right|\right]$
$\leq\mathbb{E}\left[\mathbb{E}_{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\left|f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\right|\right]\right]$
$=\delta.$
The approximation error is bounded as follows,
$\quad\mathbb{E}\left[\left|\mathbb{E}_{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|\right]$ (B10)
$=\mathbb{E}\left[\left|\int f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}-\int f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}\right|\right]$
$=\mathbb{E}\left[\left|\int f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\frac{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}\right.\right.$
$\quad-\left.\left.\int f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}\right|\right]$
$=\mathbb{E}\left[\left|\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\left(f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\frac{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\right)\right]\right|\right]$
$\leq\mathbb{E}\left[\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\left|f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\frac{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\right|\right]\right]$
$\leq\mathbb{E}\left[\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\left|\left(\frac{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-1\right)f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\right.\right.\right.$
$\quad\left.\left.\left.+f^{\prime}(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})(\mathbf{x}_{S}-\mathbf{x}_{S}^{t})+o(\mathbf{x}_{S}-\mathbf{x}_{S}^{t})\right|\right]\right]$
$\leq\epsilon_{1}f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})+\epsilon_{1}^{\prime}$
$\leq\epsilon_{2}.$
By combining Eqs. (B8), (B9), and (B10), the error bound of VAEAC approach can be obtained.
$\small\mathbb{E}[|\hat{v}(S)-v(S)|]=\frac{\sqrt{\pi}}{\sqrt{2K}}+\epsilon_{2}+\delta.$ (B11)
It can be seen from Eq.(B11), comparing with EmSHAP, the error bound of VAEAC contains the term $\delta$ that is related to the amortization error.